### v2

# Weyl-invariants of totally disconnected locally compact groups acting cocompactly on buildings

Abstract.
In several instances, the invariants of compactly generated totally disconnected locally compact groups acting on locally finite buildings can be conveniently described via invariants of the Coxeter group representing the type of the building. For certain totally disconnected locally compact groups acting on buildings, we establish and collect several results concerning, for example, the rational discrete cohomological dimension (cf. Thm. A), the flat-rank (cf. Thm. C) and the number of ends (cf. Cor. K). Moreover, for an arbitrary compactly generated totally disconnected locally compact group, we express the number of ends in terms of its cohomology groups (cf. Thm. J). Furthermore, generalising a result of F. Haglund and F. Paulin, we prove that visual graph of groups decompositions of a Coxeter group $(W,S)$ can be used to construct trees from buildings of type $(W,S)$. We exploit the latter result to show that all $\sigma$-compact totally disconnected locally compact groups acting chamber-transitively on buildings can be decomposed accordingly to any visual graph of groups decomposition of the type $(W,S)$ (cf. Thm. F and Cor. G).
Key words and phrases:
Coxeter groups, buildings, t.d.l.c. groups, invariants, ends, cohomological dimension, algebraic rank, flat rank
2010 Mathematics Subject Classification:
22D05, 20J05, 20J06
1. Introduction
It seems conceivable that certain invariants and/or properties of compactly generated, totally disconnected locally compact (= t.d.l.c.) groups acting cocompactly on a locally finite building of type $(W,S)$ are given in terms of suitable invariants or properties of $(W,S)$. The purpose of this paper is to establish some concrete new results that fit in this general framework as well as extending some classical constructions. The standard examples of t.d.l.c. groups to which these results apply are the complete Kac–Moody groups defined over a finite field (cf. [43]) and the semisimple algebraic groups defined over a non-Archimedean local field (cf. [10, 11]). However, our results also apply to more recently established classes of t.d.l.c. groups acting on buildings such as the automorphism groups of locally finite right-angled buildings and certain universal subgroups (cf. [12, 25]). Several results of this paper will concern properties and invariants of a t.d.l.c. group $G$ acting on a locally finite building of type $(W,S)$ with compact open stabilisers. In this case, a property $\mathscr{P}$ of $G$ will be called a Weyl-property if, whenever $W$ has property $\mathscr{P}$, then $G$ has property $\mathscr{P}$. Similarly, a Weyl-invariant is an invariant of $G$ which coincides with the analogous invariant of $W$.
Throughout the paper, a building $\Delta$ consists of a set $\mathcal{C}$ of chambers, together with a $W$-valued distance function $\delta\colon\mathcal{C}\times\mathcal{C}\to W$ (where $(W,S)$ is a Coxeter group with fundamental set of generators $S$) and a family of equivalence relations on $\mathcal{C}$ indexed by $S$. Indeed, associated to any building, there is a Coxeter group $(W,S)$ called the type of $\Delta$ (cf. [1, §4.1]). A building $\Delta$ is locally finite if |$|\{d\in\mathcal{C}\mid\delta(c,d)=s\}|$ is finite for all $c\in\mathcal{C}$ and $s\in S$. We will make extensive use of the so-called Davis’ realisation $|\Delta_{\mathrm{Dav}}|$ of $\Delta$ (cf. Section 2.5). It is a topological space whose compactly supported cohomology $\mathrm{H}_{c}^{*}(|\Delta_{\mathrm{Dav}}|,\mathbb{Z})$ has been explicitly computed by M. W. Davis, J. Dymara, T. Januszkiewicz, J. Meier and B. Okun in [23]. These explicit formulae fit perfectly for our need thanks to the following result.
Theorem A (cf. Lemma 3.2 and Theorem 3.4).
Let $G$ be a t.d.l.c. group acting on a locally finite building $\Delta$ of type $(W,S)$ with compact open stabilisers and finitely many orbits. Then, for every $k\geq 0$, one has a canonical isomorphism of $\mathbb{Q}$-vector spaces
(1.1) $\mathrm{dH}^{k}(G,C_{c}(G,\mathbb{Q}))\simeq\mathrm{H}_{c}^{k}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q}),$
where $\mathrm{dH}^{k}(G,C_{c}(G,\mathbb{Q}))$ denotes the rational discrete $k^{th}$-cohomology group of $G$, as introduced in [19], with coefficients in the space $C_{c}(G,\mathbb{Q})$ of continuous functions with compact support from $G$ to the discrete space $\mathbb{Q}$.
Moreover, if $\mathrm{cd}_{\mathbb{Q}}(G)$ denotes the rational discrete cohomological dimension of $G$ as defined in [19, §3.4], then $\mathrm{cd}_{\mathbb{Q}}(G)=\mathrm{cd}_{\mathbb{Q}}(W).$ In particular $G$ is compact if, and only if, $W$ is finite.
Theorem A shows in particular that the rational discrete cohomological dimension is a Weyl-invariant of $G$ for cocompact actions on $\Delta$.
Remark 1.1.
Apart from the discrete case (cf. [47]), to the authors’ knowledge it is currently unknown whether the rational discrete cohomological dimension of a compactly generated t.d.l.c. group is a quasi-isometric invariant.
If $\mathrm{cd}_{\mathbb{Q}}(W)=1$, i.e., $W$ is infinite and virtually free, Theorem A implies that $\mathrm{cd}_{\mathbb{Q}}(G)=1$ and, therefore, $\mathrm{cd}_{\mathbb{Q}}(H)\leq 1$ for every closed subgroup $H$ of $G$ (cf. [19, Proposition 3.7]). In particular, in this case, every discrete subgroup $D$ of $G$ satisfies $\mathrm{cd}_{\mathbb{Q}}(D)\leq 1$ and, consequently, every torsion-free discrete subgroup of $G$ must be free (cf. [27, Theorem 1.1] and [49, §I.3.3, Theorem 4]). Hence, Theorem A can be regarded as a generalised version of Ihara’s theorem (cf. [32]) and yields the following consequence.
Corollary B (cf. Corollary 3.5).
Let $G$ be a unimodular t.d.l.c. group acting on a locally finite building of type $(W,S)$ with compact open stabilisers and finitely many orbits. If $W$ is infinite and virtually free, then $G$ is non-compact and has a uniform lattice which is free and finitely generated. In particular, having a cocompact discrete free subgroup is a Weyl-property of $G$ for cocompact actions.
For t.d.l.c. groups acting on buildings Theorem A provides an efficient method for computing the rational discrete cohomological dimension by taking advantage of the combinatorics of Coxeter groups. E.g., the rational cohomological dimension of a Coxeter group can be computed by the Bestvina’s method (cf. [4]). In the literature there are already several results relating invariants of a t.d.l.c. group $G$ acting with compact open stabilisers and finitely many orbits on a building of type $(W,S)$ to invariants of the Coxeter group $(W,S)$. E.g., in [14, pp. 23-24] the authors show that $G$ has one end if, and only if, $W$ has one end. This can be also deduced by combining Theorem A and Theorem J, as shown in Proposition 7.7. In [3, 13], for sufficiently transitive actions, the authors relate the flat rank of $G$ to the algebraic rank of $W$, namely $\mathrm{flat}\text{-}\mathrm{rk}(G)\leq\mathrm{alg}\text{-}\mathrm{rk}(W)$. The equality holds for complete Kac–Moody groups defined over a finite field and semisimple algebraic groups over a non-Archimedean local field. Since $\mathrm{alg}\text{-}\mathrm{rk}(W)\leq\mathrm{cd}_{\mathbb{Q}}(W)$, Theorem A implies the following.
Theorem C (cf. Theorem 4.2).
Let $\Delta$ be a locally finite building and $G$ a Weyl-transitive closed subgroup of $\mathrm{Aut}_{0}(\Delta)$. Then
(1.2) $\mathrm{flat}\text{-}\mathrm{rk}(G)\leq\mathrm{cd}_{\mathbb{Q}}(G).$
For arbitrary t.d.l.c. groups the inequality (1.2) is known to hold in the case that the t.d.l.c. group $G$ is $N$-compact, i.e., every compact open subgroup of $G$ has compact normalizer (cf. [19, Proposition 3.10]). A t.d.l.c. group acting Weyl-transitively on a locally finite thick building $\Delta$ of type $(W,S)$ is $N$-compact. This result is proved in Proposition 4.1 and relies on the double-coset property. Provided $\Delta$ is thick, this fact gives an alternative proof of Theorem C.
Question.
For which non-discrete t.d.l.c. groups does $\mathrm{flat}\text{-}\mathrm{rk}(G)=\mathrm{cd}_{\mathbb{Q}}(G)$ hold?
Clearly, $\mathrm{flat}\text{-}\mathrm{rk}(G)=0=\mathrm{cd}_{\mathbb{Q}}(G)$ whenever $G$ is compact. Therefore, we are interested in non-compact examples. For complete Kac–Moody groups over a finite field and semisimple algebraic groups over a non-Archimedean local field the equality $\mathrm{flat}\text{-}\mathrm{rk}(G)=\mathrm{alg}\text{-}\mathrm{rk}(W)$ holds (cf. [3, Theorem 17] and [13, Corollary C]). Hence the answer to the question above for the latter classes of groups is provided by a comparison of the invariants $\mathrm{alg}\text{-}\mathrm{rk}(W)$ and $\mathrm{cd}_{\mathbb{Q}}(W)$. In [19], it has been already pointed out that
$\mathrm{flat}\text{-}\mathrm{rk}(G)=\mathrm{alg}\text{-}\mathrm{rk}(W)=\mathrm{cd}_{\mathbb{Q}}(W)=\mathrm{cd}_{\mathbb{Q}}(G)$
whenever $G$ is semisimple algebraic over a non-Archimedean local field. For hyperbolic Coxeter groups we will prove the following.
Theorem D (cf. Propositions 4.7 and 4.9).
For a hyperbolic Coxeter group $(W,S)$ the following hold:
  1. (a)
    If $(W,S)$ is compact, then $\mathrm{alg}\text{-}\mathrm{rk}(W)=1$ and $\mathrm{cd}_{\mathbb{Q}}(W)=|S|-1\geq 2$.
  2. (b)
    If $(W,S)$ is non-compact, then $\mathrm{alg}\text{-}\mathrm{rk}(W)=\mathrm{cd}_{\mathbb{Q}}(W)=|S|-2.$
Thanks to a celebrated decomposition theorem (cf. [26, Theorem 6.10], [52, 51]) established by J. R. Stallings in the late 60s, one knows that a finitely generated group with more than one end splits non-trivially over a finite subgroup either as amalgamated free product or as HNN-extension. If $(W,S)$ is a Coxeter group with more than one end, then $W$ is generated by involutions and thus every group homomorphism $\phi\colon W\to\mathbb{Z}$ is trivial. Hence $W$ necessarily decomposes as an amalgamated free product. Two subsets $S_{\blacktriangledown},S_{\blacktriangle}\subseteq S$, with $S_{\bullet}:=S_{\blacktriangledown}\cap S_{\blacktriangle}$, satisfying
  • (i)
    $S=S_{\blacktriangledown}\cup S_{\blacktriangle}$; and,
  • (ii)
    $m_{st}=\infty$ for all $s\in S_{\triangledown}:=S_{\blacktriangledown}\setminus S_{\bullet}$ and $t\in S_{\vartriangle}:=S_{\blacktriangle}\setminus S_{\bullet}$;
[…]
Theorem E (cf. Theorem 6.2).
For a Coxeter group $(W,S)$ the following properties are equivalent:
  • (i)
    $W$ has more than one end;
  • (ii)
    $(W,S)$ admits a non-trivial spherical $\infty$-decomposition $S=S_{\blacktriangledown}\cup S_{\blacktriangle}$, where being non-trivial means that $S_{\triangledown}$ and $S_{\vartriangle}$ are non-empty, and spherical that $(W_{S_{\bullet}},S_{\bullet})$ is a finite Coxeter group.
Remark 1.2.
The notion of $\infty$-decomposition has been considered in [40] with the name of “visual splitting” and in [30] with the name of “scindement”.
Example 1.1.
If the Coxeter matrix of $(W,S)$ has only finite entries, from Theorem E one concludes that $W$ is either finite or $1$-ended. For instance, if $(W,S)$ is a hyperbolic Coxeter group of rank at least $4$, then $W$ has one end (cf. [31, Figures 6.2 and 6.3]). If $(W,S)$ is affine, then $W$ has one or two ends, and it has two ends if, and only if, $W\simeq D_{\infty}$ (cf. [31, Figure 4.1]).
Let $\Delta$ be a building of type $(W,S)$. In [30], F. Haglund and F. Paulin showed that if $(W,S)$ admits an $\infty$-decomposition $S=S_{\blacktriangledown}\cup S_{\blacktriangle}$ then $\Delta$ can be used to construct a graph $\Gamma_{\blacktriangledown\!/\!\blacktriangle}(\Delta)$ which turns out to be a tree. More precisely, given $S_{\blacktriangledown},S_{\blacktriangle}\subseteq S$ such that $S=S_{\blacktriangledown}\cup S_{\blacktriangle}$ and $S_{\bullet}=S_{\blacktriangledown}\cap S_{\blacktriangle}$, the following are known to be equivalent:
  • (i)
    $S=S_{\blacktriangledown}\cup S_{\blacktriangle}$ is an $\infty$-decomposition of $(W,S)$;
  • (ii)
    $W\simeq W_{S_{\blacktriangledown}}\amalg_{W_{S_{\bullet}}}W_{S_{\blacktriangle}}$;
  • (iii)
    $\Gamma_{\blacktriangledown\!/\!\blacktriangle}(\Delta)$ is a tree.
[…]
Theorem F (cf. Theorem 5.7).
Let $\Delta=(\mathcal{C},\delta)$ be a building of type $(W,S)$ and $(\mathcal{G},\Lambda)$ be a graph of special subgroups of $(W,S)$. Then the following conditions are equivalent:
  • (i)
    $(\mathcal{G},\Lambda)$ is a visual graph of groups decomposition of $(W,S)$;
  • (ii)
    $S=\bigcup_{v\in\mathscr{V}(\Lambda)}S_{v}$ and $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C})$ is a tree.
For a finite tree $\Lambda$ one says that the (topological) group $G$ has a $\Lambda$-tree action, if there exists a tree $\mathscr{T}$ with a (discrete) left $G$-action without inversion of edges satisfying $G\backslash\mathscr{T}\simeq\Lambda$. Theorem F has the following consequence showing that having a $\Lambda$-tree action is a Weyl-property of a $\sigma$-compact t.d.l.c. group for chamber-transitive actions on $\Delta$.
Corollary G (cf. Corollary 5.9).
Let $G$ be a $\sigma$-compact t.d.l.c. group acting chamber-transitively on a building $\Delta=(\mathcal{C},\delta)$ of type $(W,S)$, and let $c\in\mathcal{C}$. Suppose that $(W,S)$ admits a visual graph of groups decomposition $(\mathcal{G},\Lambda)$. Then $G$ is topologically isomorphic to the fundamental group of the tree of t.d.l.c. groups $(\widetilde{\mathcal{G}},\Lambda)$, where $\widetilde{\mathcal{G}}_{\lambda}$ is the setwise stabiliser of $\mathrm{Res}_{S_{\lambda}}(c)$ for every $\lambda\in\mathscr{V}(\Lambda)\sqcup\mathscr{E}(\Lambda)$, and the continuous open monomorphisms $\widetilde{\mathcal{G}}_{e}\hookrightarrow\widetilde{\mathcal{G}}_{t(e)}$, $e\in\mathscr{E}(\Lambda)$, are given by inclusion.
Moreover, the universal Bass–Serre tree of ${\pi_{1}(\widetilde{\mathcal{G}},\Lambda)}$ is isomorphic to $\Gamma_{(\widetilde{\mathcal{G}},\Lambda)}^{\Delta}(\mathcal{C})$.
By a repeated application of Theorem E, M. W. Davis observed that every Coxeter group $(W,S)$ is accessible, i.e., $(W,S)$ admits a visual graph of groups decomposition $(\mathcal{G},\mathscr{T})$ in which each vertex-group is at most $1$-ended and each edge-group is finite. If, in addition, $W$ is infinite but every vertex-group of $(\mathcal{G},\mathscr{T})$ is a spherical special subgroup of $W$, we say that $\pi_{1}(\mathcal{G},\mathscr{T})$ is a completely spherical visual graph of groups decomposition of $(W,S)$. This property can be characterised as follows.
Theorem H (cf. Theorem 6.4).
Let $(W,S)$ be a Coxeter group. Then the following are equivalent:
  • (i)
    $W$ is infinite and virtually free;
  • (ii)
    $\mathrm{vcd}(W)=1$ (or, equivalently, $\mathrm{cd}_{\mathbb{Q}}(W)=1$);
  • (iii)
    $(W,S)$ admits a completely spherical $\infty$-decomposition;
  • (iv)
    $(W,S)$ is not spherical, but every clique $\Lambda\subseteq\Gamma_{\infty}(W,S)$; is spherical, and $\Gamma_{\infty}(W,S)$ is chordal.
The equivalence of (iii) and (iv) was already shown in [40]. Alternatively, one may deduce this result from the following more general decomposition theorem, which deals with visual graph of groups decompositions whose groups satisfy a given property $\mathscr{P}$.
Theorem I (cf. Theorem 6.6).
Let $(W,S)$ be an infinite Coxeter group such that $\Gamma_{\infty}(W,S)$ has the $\mathscr{P}$-clique property (cf. Subsection 6.4). If $\Gamma_{\infty}(W,S)$ is chordal, then $W$ admits a completely-$\mathscr{P}$infty-decomposition $\pi_{1}(\mathcal{G},\mathscr{T},x_{0})$ (i.e., one has an isomorphism $W\simeq\pi_{1}(\mathcal{G},\mathscr{T},x_{0})$ where $(\mathcal{G},\mathscr{T})$ is a graph of special subgroups such that all vertex-groups and edge-groups satisfy property $\mathscr{P}$).
In geometric group theory it is well-known that, given a discrete group $\Gamma$, the cohomology of $\Gamma$ with coefficients in $\mathbb{Z}\Gamma$ is isomorphic to the compactly supported cohomology of $Y$, where $Y$ is any locally finite contractible CW-complex on which $\Gamma$ acts properly and cocompactly (see for example [28, Part IV]). By [19, Section 6.8], this result carries over to the realm of t.d.l.c. groups and the rational discrete cohomology. In particular, if $G$ is a t.d.l.c. group acting cocompactly on a contractible CW-complex $X$ with compact open stabilisers, then the cohomology with compact support of $X$ can be used to compute the rational discrete cohomological dimension of $G$ and determine whether $G$ is a rational discrete duality t.d.l.c. group (cf. [19, Theorem 6.7]). For a discrete group $\Gamma$, the compactly supported cohomology of $Y$ also determines the number of ends of $\Gamma$. The t.d.l.c. analogue can be obtained as a consequence of the following result and Remark 3.3.
Theorem J (cf. Theorem 7.3).
Let $G$ be a compactly generated t.d.l.c. group. Then
$e(G)=1-\dim_{\mathbb{Q}}\mathrm{dH}^{0}(G,\mathrm{\bf Bi}(G))+\dim_{\mathbb{Q}}\mathrm{dH}^{1}(G,\mathrm{\bf Bi}(G)),$
where $\mathrm{\bf Bi}(G)$ is the rational $\mathbb{Q}[G]$-bimodule introduced in [19] as a suitable substitute for the group algebra (which is non-canonically isomorphic to $C_{c}(G,\mathbb{Q})$).
Finally, we investigate some consequences of Theorem H for t.d.l.c. groups acting cocompactly on locally finite buildings with compact open stabilisers. We exploit it to provide an alternative proof to the fact that $G$ has one end if, and only if, $W$ has one end (cf. Proposition 7.7). In particular we obtain the following:
Corollary K (cf. Remark 7.6, Proposition 7.7, Remark 7.8).
Being 00- or $1$-ended is a Weyl property of $G$ for cocompact actions. However, the number of ends is not a Weyl invariant.
Conventions and notations
Throughout the paper $\sqcup$” will denote the disjoint union operator of sets. For a Coxeter group $(W,S)$ the set $S$ will always be finite. Every affine (as well as hyperbolic) Coxeter group is supposed to be irreducible. Whenever a topological group $G$ acts on a graph, we assume that $G$ has open vertex stabilisers. In order to avoid an overload of notation for a graph of groups $(\mathcal{G},\Lambda)$ we will denote by $\pi_{1}(\mathcal{G},\Lambda)$ the fundamental group of $(\mathcal{G},\Lambda)$ omitting the choice of the base point.
Acknowledgements
The authors would like to thank P.-E. Caprace for his useful comments. The first and second named authors were supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB-TRR 358/1 2023 – 491392403. The third author gratefully acknowledges financial support by the PRIN2022 “Group theory and its applications”. All authors are members of the Gruppo Nazionale per le Strutture Algebriche, Geometriche e le loro Applicazioni (GNSAGA), which is part of the Istituto Nazionale di Alta Matematica (INdAM).
2. Preliminaries
2.1. Graphs
An undirected graph $(V,E)$ consists of a non-empty set of vertices $V$ together with a set of edges $E\subseteq\mathscr{P}_{2}(V)$, where $\mathscr{P}_{2}(V)$ denotes the set of all subsets of $V$ of cardinality $2$. A subgraph $(U,F)$ of $(V,E)$ is an undirected graph such that $U\subseteq V$ and $F\subseteq E$. An undirected graph $(V,E)$ is said to be complete if $E=\mathscr{P}_{2}(V)$. A complete subgraph $(U,F)$ of an undirected graph $(V,E)$ satisfying $|U|=n$ is said to be an $n$-clique. A subgraph $(U,F)$ of an undirected graph $(V,E)$ is said to be induced if $F=\mathscr{P}_{2}(U)\cap E$. For instance, any $n$-clique is an induced subgraph. A path in $(V,E)$ is a finite sequence of vertices $(v_{0},\ldots,v_{n})$ such that $\{v_{i},v_{i+1}\}\in E$ for every $10\leq i\leq n-$ 1.
Every undirected graph $\Gamma=(V,E)$ defines a combinatorial graph $\vec{\Gamma}=(V,\vec{E})$ in the sense of J-P. Serre (cf. [49, §I.2.1, Definitions 1 and 5]), where
(2.1) $\vec{E}=\{(v,w),(w,v)\mid\{v,w\}\in E\},$
the origin map $o\colon\vec{E}\to V$ is the projection on the first coordinate, the terminus map $t\colon\vec{E}\to V$ is the projection on the second coordinate, and the edge-inversion map ¯$\bar{\cdot}\colon\vec{E}\to\vec{E}$ interchanges the first and second coordinates. On the other hand, every combinatorial graph $\Gamma=(\mathscr{V},\mathscr{E})$ in the sense of J-P. Serre defines an undirected graph $\ddot{\Gamma}=(\mathscr{V},\ddot{\mathscr{E}})$ given by
(2.2) ¨=$\ddot{\mathscr{E}}=\big{\{}\{o(\mathbf{e}),t(\mathbf{e})\}\mid\mathbf{e}\in\mathscr{E}\big{\}}.$
2.2. Coxeter groups and associated diagrams
A Coxeter matrix $[m_{st}]_{s,t\in S}$ is a symmetric matrix with diagonal entries equal to $1$ and with non-diagonal entries contained in $\mathbb{Z}_{\geq 2}\cup\{\infty\}$. One can associate to $M$ the Coxeter group $(W,S)$, which is the group generated by $S$ subject to the relations $(st)^{m_{st}}=1$, whenever $m_{st}\neq\infty$. It turns out that $m_{s,t}$ coincides with the order of $st$ in $W$. A special subgroup $W_{J}$ of $(W,S)$ is a subgroup generated by a subset $J$ of $S$. Notice that $(W_{J},J)$ is a Coxeter group for every $J\subseteq S$ (cf. [31, Theorem 5.5(a)]). A Coxeter group $(W,S)$ is said to be spherical if $W$ is finite. One says that $J\subseteq S$ is spherical if $(W_{J},J)$ is spherical. For further details the reader is referred to [31].
Classically, to each Coxeter group $(W,S)$ one associates the so-called Coxeter diagram $\Gamma(W,S)$, which is the edge-labelled undirected graph with set of vertices $S$, set of edges $E(\Gamma(W,S))=\big{\{}\{s,t\}\mid s,t\in S,\,m_{st}\neq 2\big{\}}$ and edge-labelling $m\colon E(\Gamma(W,S))\to\mathbb{Z}_{\geq 3}\cup\{\infty\}$, $m(\{s,t\}):=m_{st}$. Usually, one omits the label of all edges $\{s,t\}$ in $\Gamma(W,S)$ satisfying $m_{st}=3$. Sometimes, it is convenient to consider another graph $\Gamma_{\infty}(W,S)$ associated to $(W,S)$, the presentation diagram of $(W,S)$ (cf. [40]). It is the edge-labelled undirected graph with vertex set $S$ and edge set $E(\Gamma_{\infty}(W,S))=\big{\{}\{s,t\}\mid s,t\in S,\,m_{st}\neq\infty\big{\}}$, in which every edge $\{s,t\}$ is labelled with $m_{st}$. Each of the two edge-labelled undirected graphs, $\Gamma(W,S)$ and $\Gamma_{\infty}(W,S)$, determines uniquely the Coxeter matrix $[m_{st}]_{s,t\in S}$ and thus the Coxeter group $(W,S)$.
2.3. Virtual cohomological dimension of Coxeter groups
Let $\Gamma$ be a virtually torsion-free group. A result due to J-P. Serre (cf. [48]) shows that all finite-index torsion-free subgroups of $\Gamma$ have the same cohomological dimension over a non-zero associative ring $R$, which is then defined to be the virtual cohomological dimension $\mathrm{vcd}_{R}(\Gamma)$ of $\Gamma$ over $R$. From now on $\mathrm{vcd}_{\mathbb{Z}}$ will be denoted by $\mathrm{vcd}$. In 1993, M. Bestvina (cf. [4]) provided an effective method to compute $\mathrm{vcd}_{R}(W)$ of a finitely generated Coxeter group $(W,S)$, where $R$ is either the integers or a prime field. The construction, recalled below, produces an $R$-acyclic complex $B_{R}(W,S)$ of dimension $\mathrm{vcd}_{R}(W)$.
A subset $X\subseteq\mathbb{R}^{n}$ is a polyhedron if each point $a\in X$ has a cone neighbourhood $N=aL$ in $X$ with $L$ compact (cf. [46]). Recall that a polyhedron can always be triangulated as a locally finite simplicial complex. Let $(W,S)$ be a finitely generated Coxeter group and denote by $\mathfrak{S}$ the family of all spherical subsets of $S$ ordered by inclusion. For every maximal element $J\in\mathfrak{S}$ define $P_{F}$ to be a point. Let $J\in\mathfrak{S}$ be non-maximal and assume that $P_{J^{\prime}}$ has been constructed for all $J^{\prime}\in\mathfrak{S}$ such that $J^{\prime}\supset J$. Define $P_{J}$ to be an $R$-acyclic polyhedron containing $\bigcup_{{}_{J^{\prime}\supset J}}P_{J^{\prime}}$ of the least possible dimension. The construction ends once $P_{\emptyset}$ has been built.
In Section 4 the Bestvina’s complex $B_{R}(W,S)$ is used to compute the rational cohomological dimension of hyperbolic Coxeter groups. Moreover, for arbitrary Coxeter groups, one can use the Bestvina’s complex (and the associated augmented cellular complex) to deduce that $\mathrm{cd}_{\mathbb{Q}}(W)\leq\mathrm{vcd}(W)<\infty.$ Here $\mathrm{cd}_{\mathbb{Q}}(W)$ denotes the rational cohomological dimension of $W$, i.e.,
(2.3) $\mathrm{cd}_{\mathbb{Q}}(W)=\max\{n\geq 0\mid\mathrm{H}^{n}(W,\mathbb{Q}[W])\neq 0\}$
(cf. [9, Chapter VIII, Proposition 6.7]).
2.4. Buildings
An example of a building of type $(W,S)$ is the abstract Coxeter complex $\Sigma(W,S)=(W,\delta_{W})$ of type $(W,S)$, where $\delta_{W}(w_{1},w_{2})=w_{1}^{-1}w_{2}$ for all $w_{1},w_{2}\in W$ (cf. [22, Example 18.1.3]). Given two buildings $\Delta=(\mathcal{C},\delta)$ and $\Delta^{\prime}=(\mathcal{C}^{\prime},\delta^{\prime})$ of type $(W,S)$, a map $\varphi\colon\mathcal{C}\to\mathcal{C}^{\prime}$ is a $W$-isometry if, for all $c_{1},c_{2}\in\mathcal{C}$,
(2.4) $\delta^{\prime}(\varphi(c_{1}),\varphi(c_{2}))=\delta(c_{1},c_{2}).$
Such a map is necessarily injective. Following [22, p. 333], for every building $\Delta=(\mathcal{C},\delta)$ of type $(W,S)$ there is a unique $W$-isometry $W\to\mathcal{C}$ mapping $1_{W}$ to a prescribed chamber $c\in\mathcal{C}$. An apartment of $\Delta$ is any $W$-isometric image of $W$ in $\mathcal{C}$. The following fact collects some properties about apartments in a building.
Fact 2.1 (cf. [22, p. 333] and [36, §4]).
Let $\Delta=(\mathcal{C},\delta)$ be a building of type $(W,S)$.
  • (i)
    For every apartment $\Sigma$ of $\Delta$, the chamber system $(\mathcal{C}(\Sigma),\delta|_{\mathcal{C}(\Sigma)})$ is a building of type $(W,S)$.
  • (ii)
    There exists a collection of apartments $\mathfrak{A}$, called atlas of $\Delta$, with the following property: every two chambers $c$ and $d$ of $\Delta$ are both contained in the set of chambers of some $\Sigma\in\mathfrak{A}$. In particular, one has
    $\mathcal{C}=\bigcup_{\Sigma\in\mathfrak{A}}\mathcal{C}(\Sigma).$
A gallery $\gamma$ in $\Delta$ is a sequence of chambers $(c_{0},c_{1},\ldots,c_{n})$ satisfying $\delta(c_{i},c_{i+1})=s_{i}\in S$ for every $10\leq i\leq n-$ 1. The sequence $(s_{1},\ldots,s_{n})$ is called the type of $\gamma$. For all $J\subseteq S$ and $c\in\mathcal{C}$, the $J$-residue of $\Delta$ centred at $c$ is defined as
$\mathrm{Res}_{J}(c):=\{d\in\mathcal{C}\mid\delta(c,d)\in W_{J}\}.$
In other words, $\mathrm{Res}_{J}(c)$ is the collection of all $d\in\mathcal{C}$ for which either $d=c$ or there is a gallery $(c_{0}=c,c_{1},\ldots,c_{n}=d)$ in $\Delta$ whose type $(s_{1},\ldots,s_{n})$ satisfies $s_{i}\in J$ for every $1\leq i\leq n$ (cf. [1, Definition 5.26]). In particular, for all $c,d\in\mathcal{C}$ and $J\subseteq S$, one deduces that
(2.5) $\mathrm{Res}_{J}(c)=\mathrm{Res}_{J}(d)\Longleftrightarrow\delta(c,d)\in W_{J}.$
A $J$-residue is called spherical if $J$ is a spherical subset of $S$. The building $\Delta$ is said to be locally finite if, for every $s\in S$, all $\{s\}$-residues are finite. Denote by $\mathcal{R}$ the set of all residues of $\Delta$, by $\mathcal{R}_{\mathrm{sph}}$ the set of all spherical ones and, given $J\subseteq S$, let $\mathcal{R}_{J}$ be the set of all $J$-residues of $\Delta$.
2.5. The Davis’ complex of a building
Following [36, §5], we recall here a notable simplicial complex associated to a building $\Delta=(\mathcal{C},\delta)$. The Davis’ complex $\Delta_{\mathrm{Dav}}$ of $\Delta$ is the simplicial complex whose $k$-simplices, for $k\geq 0$, are all chains
(2.6) $\mathrm{Res}_{J_{0}}(c)\subsetneq\cdots\subsetneq\mathrm{Res}_{J_{k}}(c),$
where $c\in\mathcal{C}$ and $\mathrm{Res}_{J_{0}}(c),\dots,\mathrm{Res}_{J_{k}}(c)\in\mathcal{R}_{\mathrm{sph}}$ with $J_{0}\subsetneq J_{1}\subsetneq\cdots\subsetneq J_{m}$. The simplicial complex $\Delta_{\mathrm{Dav}}$ is locally finite if, and only if, $\Delta$ is locally finite (cf. [36, p. 22]). By [22, Corollary 18.3.6], the geometric realisation $|\Delta_{\mathrm{Dav}}|$ of $\Delta_{\mathrm{Dav}}$ is contractible (with respect to the weak topology) and is usually called the Davis’ realisation of $\Delta$. Moreover, the augmented cellular chain complex associated to $|\Delta_{\mathrm{Dav}}|$ is exact (cf. [19, §A.3]).
2.6. Groups acting on buildings
Given a building $\Delta=(\mathcal{C},\delta)$ of type $(W,S)$, its group of type-preserving automorphisms is defined as
(2.7) $\mathrm{Aut}_{0}(\Delta):=\{g\mid g:\mathcal{C}\to\mathcal{C}\text{ surjective}W\text{-isometry of }\Delta\}.$
The group $\mathrm{Aut}_{0}(\Delta)$ is a totally disconnected group with respect to the permutation topology, i.e., the group topology having the family $\{\mathrm{stab}_{\mathrm{Aut}_{0}(\Delta)}(F)\mid F\subseteq\mathcal{C}\text{ finite}\}$ as a neighbourhood basis of the identity. Here $\mathrm{stab}_{\mathrm{Aut}_{0}(\Delta)}(F)$ denotes the pointwise stabiliser of the set $F\subseteq\mathcal{C}$. If $\Delta$ is locally finite, $\mathrm{stab}_{\mathrm{Aut}_{0}(\Delta)}(F)$ is profinite for every finite subset $F$ of $\mathcal{C}$, and therefore $\mathrm{Aut}_{0}(\Delta)$ is locally compact (cf. [36, Theorem 5.1]).
A topological group $G$ acts on $\Delta$ if there is a continuous group homomorphism $\varphi\colon G\to\mathrm{Aut}_{0}(\Delta)$. For all $g\in G$ and $c\in\mathcal{C}$, one writes $g\cdot c$ in place of $\varphi(g)(c)$. A group $G$ acts Weyl-transitively on $\Delta$ if, for every $w\in W$, the $G$-action is transitive on the set
(2.8) $\{(c,d)\in\mathcal{C}\times\mathcal{C}\mid\delta(c,d)=w\}.$
3. The rational discrete cohomological dimension of a t.d.l.c. group acting on a building
Given a locally finite building $\Delta$, denote by $\mathrm{H}_{c}^{\bullet}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q})=\mathrm{H}_{c}^{\bullet}(|\Delta_{\mathrm{Dav}}|,\mathbb{Z})\otimes_{\mathbb{Z}}\mathbb{Q}$ the cohomology with compact support of the geometric realisation of $\Delta_{\mathrm{Dav}}$ with rational coefficients. Furthermore, let $C_{c}(G,\mathbb{Q})$ be the space of all locally constant functions from $G$ to $\mathbb{Q}$ with compact support and let $\mathrm{cd}_{\mathbb{Q}}(W)$ be the ordinary rational cohomological dimension of $W$ (cf. Section 2.3). Moreover, for a t.d.l.c. group $G$ we denote by $\mathrm{cd}_{\mathbb{Q}}(G)$ the rational discrete cohomological dimension of $G$ as defined in [19, §3.4].
Lemma 3.1.
For every locally finite building $\Delta$ of type $(W,S)$, one has
$\mathrm{cd}_{\mathbb{Q}}(W)=\max\{n\geq 0:\mathrm{H}_{c}^{n}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q})\neq 0\}.$
Proof.
The argument used to prove the equality (6.15) in [19] can be transferred verbatim. ∎
Lemma 3.2.
Let $G$ be a t.d.l.c. group acting on a locally finite building $\Delta$ of type $(W,S)$ with compact stabilisers and finitely many orbits. Then, for every $k\geq 0$, one has a canonical isomorphism of $\mathbb{Q}$-vector spaces
(3.1) $\mathrm{dH}^{k}(G,C_{c}(G,\mathbb{Q}))\simeq\mathrm{H}_{c}^{k}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q}).$
Proof.
The simplicial complex $\Delta_{\mathrm{Dav}}$ is finite-dimensional, locally finite and has contractible geometric realisation (cf. Section 2.5). By hypothesis, $G$ acts on $\Delta_{\mathrm{Dav}}$ with compact open vertex stabilisers (cf. [36, Lemma 5.13]). Moreover, for all $g\in G$, $c\in\mathcal{C}$ and $J\subseteq S$, one has $\mathrm{Res}_{J}(g\cdot c)=g\cdot\mathrm{Res}_{J}(c)$ and then $G$ has finitely many orbits on each skeleton of $\Delta_{\mathrm{Dav}}$. Now the argument used to prove [19, Equation (6.13)] can be transferred verbatim. ∎
Remark 3.3.
More generally, there exist analogous isomorphisms to (3.1) when replacing $\Delta_{\mathrm{Dav}}$ by any locally finite contractible simplicial complex $X$ of type $\mathrm{F}_{\infty}$ which is acted on by $G$ with compact open stabilisers (cf. [19, Proof of Theorem 6.7 and Equation (6.13)] for details).
Theorem 3.4.
Let $G$ be a t.d.l.c. group acting on a locally finite building $\Delta$ of type $(W,S)$ with compact stabilisers and finitely many orbits. Then
(3.2) $\mathrm{cd}_{\mathbb{Q}}(G)=\mathrm{cd}_{\mathbb{Q}}(W).$
In particular $G$ is compact if, and only if, $W$ is spherical.
Proof.
Arguing as in the proof of Lemma 3.2, $G$ acts on $\Delta_{\mathrm{Dav}}$ with compact open stabilisers and finitely many orbits on each skeleton. In particular, $\mathrm{cd}_{\mathbb{Q}}(G)\leq\dim(\Delta_{\mathrm{Dav}})$ and $G$ is of type $\mathrm{FP}_{\infty}$ (cf. [19, Proposition 6.6]). By [19, Proposition 4.7, Equation (4.64)], one obtains that
$\mathrm{cd}_{\mathbb{Q}}(G)=\max\{n\geq 0\mid\mathrm{dH}^{n}(G,C_{c}(G,\mathbb{Q}))\neq 0\},$
which, by Lemma 3.2, equals $\max\{n\geq 0\mid\mathrm{H}_{c}^{n}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q})\neq 0\}.$ By Lemma 3.1, $\mathrm{cd}_{\mathbb{Q}}(G)=\mathrm{cd}_{\mathbb{Q}}(W)$. The last assertion of the statement is now a consequence of [19, Proposition 3.7(a)]. ∎
Corollary 3.5.
Let $G$ be a unimodular t.d.l.c. group acting on a locally finite building $\Delta$ of type $(W,S)$ with compact stabilisers and finitely many orbits. If $W$ is infinite and virtually free, then $G$ is non-compact and has a finitely generated free subgroup that is cocompact and discrete.
Proof.
By [27, Corollary 1.2], $\mathrm{cd}_{\mathbb{Q}}(W)=1$. Hence, by Theorem 3.4, $\mathrm{cd}_{\mathbb{Q}}(G)=1$ as well. Moreover $G$ acts on $\Delta_{\mathrm{Dav}}$ with compact open stabilisers and finitely many orbits and so $G$ is compactly presented (cf. [18, Theorem 4.9]). Thus $G$ is the fundamental group of a finite graph of profinite groups (cf. [15, Theorem B]). As $G$ is unimodular, [38, Theorem 16] yields the claim. ∎
4. A relation between the flat rank and the rational discrete cohomological dimension
This section deals with the connection of two relevant invariants associated to a t.d.l.c. group $G$, its cohomological dimension $\mathrm{cd}_{\mathbb{Q}}(G)$ and its flat rank $\mathrm{flat}\text{-}\mathrm{rk}(G)$ (cf. [3, Section 1.3]). Following [19], a t.d.l.c. group $G$ is said to be $N$-compact if, for every compact open subgroup $\mathcal{O}$ of $G$, the normalizer $N_{G}(\mathcal{O})$ is compact. By [19, Proposition 3.10(b)], for such groups one has
(4.1) $\mathrm{flat}\text{-}\mathrm{rk}(G)\leq\mathrm{cd}_{\mathbb{Q}}(G).$
The result below shows that, given a locally finite thick building $\Delta$, every t.d.l.c. group acting Weyl-transitively on $\Delta$ with compact open stabilisers is $N$-compact. As a consequence of (4.1), one deduces that for these t.d.l.c. groups one has $\mathrm{flat}\text{-}\mathrm{rk}(G)\leq\mathrm{cd}_{\mathbb{Q}}(G)$.
Proposition 4.1.
Let $G$ be a t.d.l.c. group acting Weyl-transitively on a locally finite thick building $\Delta$ with compact stabilisers. Then $G$ has the double coset property and, in particular, is $N$-compact.
Before proving the result, for the convenience of the reader we recall some terminology and notations. Let $\mathcal{O}\leq G$ be a compact open subgroup of a t.d.l.c. group. If $G$ is unimodular, one observes that
(4.2) $|N_{G}(\mathcal{O}):\mathcal{O}|=|\{\mathcal{O}g\mathcal{O}\in\mathcal{O}\backslash G/\mathcal{O}\mid\mu_{\mathcal{O}}(\mathcal{O}g\mathcal{O})=1\}|,$
where $\mu_{\mathcal{O}}$ denotes the Haar measure of $G$ such that $\mu_{\mathcal{O}}(\mathcal{O})=1$. Indeed, one has
$|\mathcal{O}:\mathcal{O}\cap g\mathcal{O}g^{-1}|=\mu_{\mathcal{O}}(\mathcal{O}g\mathcal{O})=\mu_{\mathcal{O}}(\mathcal{O}g^{-1}\mathcal{O})=|\mathcal{O}:\mathcal{O}\cap g^{-1}\mathcal{O}g|.$
In other words, for unimodular t.d.l.c. groups, the index $|N_{G}(\mathcal{O}):\mathcal{O}|$ coincides with $R_{\mathcal{O}}(1)$, where $R_{\mathcal{O}}(n):=|\{\mathcal{O}g\mathcal{O}\in\mathcal{O}\backslash G/\mathcal{O}\mid\mu_{\mathcal{O}}(\mathcal{O}g\mathcal{O})=n\}|.$ Following [17, §6], a t.d.l.c. group $G$ is said to have the double coset property if $R_{\mathcal{O}}(n)<\infty$ for every $n\geq 1$. According to [17, Proposition 6.2], this property does not depend on $\mathcal{O}$.
Proof.
By [17, Proposition 6.2], it suffices to show that $G$ has the double coset property for the stabiliser $\mathcal{O}$ of a chamber in $\Delta$. By Weyl-transitivity, there is a bijection between $\mathcal{O}\backslash G/\mathcal{O}$ and $W$ and then one may write $G=\bigsqcup_{w\in W}\mathcal{O}w\mathcal{O}$ (cf. [1, §6.1.4]). For $s\in S$, let $q_{s}=|\mathcal{O}:\mathcal{O}\cap s\mathcal{O}s^{-1}|$. By [3, Proposition 4], for every reduced word $w=s_{1}\cdots s_{n}$ in $(W,S)$ one has
$\mu_{\mathcal{O}}(\mathcal{O}w\mathcal{O})=|\mathcal{O}:\mathcal{O}\cap w\mathcal{O}w^{-1}|=q_{s_{1}}\cdots q_{s_{n}}=:q_{w}.$
Since $q_{s}\geq 2$ for every $s\in S$, there is $k\geq 1$ such that
(4.3) $\sum_{w\in W}q_{w}^{-k}\leq\sum_{w\in W}2^{-k\ell(w)}<\infty,$
where $\ell$ is the word length of $(W,S)$. Indeed, the series on the right-hand side of (4.3) is the Poincaré series $P_{(W,S)}(t)$ of $(W,S)$ evaluated in $t=2^{-k}$ which converges for all $k\gg 1$. Since $q_{w}^{-k}>0$ for all $w\in W$, for every $n\geq 1$ one deduces that
$\infty>\sum_{w\in W}q_{w}^{-k}\geq\sum_{w\in W:\atop q_{w}=n}n^{-k}=|\{w\in W\mid q_{w}=n\}|\cdot n^{-k}.$
As a conclusion, the cardinality $\mathcal{R}_{\mathcal{O}}(n)=|\{w\in W\mid q_{w}=n\}|$ is finite for every $n\geq 1$. ∎
For Weyl-transitive closed subgroups of $\mathrm{Aut}_{0}(\Delta)$, one can provide an alternative proof of (4.1) based on the relation between the invariants of $G$ and those of $W$. Recall that the algebraic rank $\mathrm{alg}\text{-}\mathrm{rk}(W)$ of $W$ is the maximal $\mathbb{Z}$-rank of free abelian subgroups of $W$. In [37, Theorem 6.8.3], D. Krammer proved that its computation can be reduced to the study of combinatorial properties of the associated Coxeter diagram.
Theorem 4.2.
Let $\Delta$ be a locally finite building and $G$ a Weyl-transitive closed subgroup of $\mathrm{Aut}_{0}(\Delta)$. Then
$\mathrm{flat}\text{-}\mathrm{rk}(G)\leq\mathrm{cd}_{\mathbb{Q}}(W)=\mathrm{cd}_{\mathbb{Q}}(G).$
Proof.
Combining [3, Theorem A] and [13, Corollary C], one deduces that
$\mathrm{flat}\text{-}\mathrm{rk}(G)\leq\mathrm{alg}\text{-}\mathrm{rk}(W).$
Then $\mathrm{alg}\text{-}\mathrm{rk}(W)\leq\mathrm{cd}_{\mathbb{Q}}(W)$ and Theorem 3.4 yields the claim. ∎
Let $\mathbb{K}$ be a non-Archimedean local field. Let $\mathbf{G}$ be a semisimple, simply connected algebraic group defined over $\mathbb{K}$ with affine Weyl group $W$. If $\mathbf{G}(\mathbb{K})$ denotes the group of $\mathbb{K}$-rational points of $\mathbf{G}$, then
(4.4) $\mathrm{alg}\text{-}\mathrm{rk}(W)=\mathrm{flat}\text{-}\mathrm{rk}(\mathbf{G}(\mathbb{K}))=\mathrm{cd}_{\mathbb{Q}}(\mathbf{G}(\mathbb{K}))=\mathrm{cd}_{\mathbb{Q}}(W)$
(cf. [3, Corollary 18] and [19, Example 3.11]).
Remark 4.3.
For $p$ prime, $(\mathrm{PSL}_{d+1}(\mathbb{Q}_{p}))_{d\geq 1}$ provides a sequence of topologically simple compactly generated t.d.l.c. groups having arbitrary rational discrete cohomological dimension $d$.
Let $G$ be a group with a locally finite twin root datum of type $(W,S)$, and denote by $(B_{\pm},N,S)$ the two associated BN-pairs (cf. [3, §1.2]). Let $\Delta_{+}$ be the locally finite building defined by $(B_{+},N,S)$, and denote by $\bar{G}$ the closure of the image of $G$ in $\mathrm{Aut}_{0}(\Delta_{+})$. The topological group $\bar{G}$ is called geometric completion of $G$ (cf. [39, p. 187]). For example, if $G$ is an abstract Kac–Moody group defined over a finite field, then $\bar{G}$ gives the so-called complete Kac–Moody group [44]. For such a group one has $\mathrm{alg}\text{-}\mathrm{rk}(W)=\mathrm{flat}\text{-}\mathrm{rk}(\bar{G})$ as well (cf. [3, Theorems A and B] and [13, Corollary C]). In particular, from Theorem 3.4 and Theorem 4.2 one deduces the following.
Proposition 4.4.
Let $G$ be a group with a locally finite twin root datum of type $(W,S)$, and denote by $\bar{G}$ the geometric completion of $G$. Then $\mathrm{alg}\text{-}\mathrm{rk}(W)=\mathrm{cd}_{\mathbb{Q}}(W)$ if, and only if, $\mathrm{flat}\text{-}\mathrm{rk}(\bar{G})=\mathrm{cd}_{\mathbb{Q}}(\bar{G})$.
Remark 4.5.
For every integer $n\geq 1$ there exists a non-linear, topologically simple, compactly generated, t.d.l.c. group $\bar{G}_{n}$ of flat rank $n$ (cf. [3, Theorem 25]). By Theorem 3.4 we can compute $\mathrm{cd}_{\mathbb{Q}}(\bar{G}_{n})$ for every $n\geq 1$ as follows. Recall that the topological Kac–Moody groups $(\bar{G}_{n})_{n\geq 1}$ are constructed by using the following sequence of connected Coxeter diagrams $(Y_{n})_{n\in\mathbb{N}}$:
    let $Y_{1}$ be a cycle of length $5$ all of whose edges are labelled $\infty$,
    if $n>1$ let $Y_{n}$ be the diagram obtained by joining every vertex of $Y_{1}$ to every vertex of a diagram of type $\tilde{A}_{n}$ by an edge labelled with $\infty$.
Denote by $(W_{n},\Sigma_{n})$ the Coxeter group associated to the Coxeter diagram $Y_{n}$. The Bestvina’s complex $B_{\mathbb{Z}}(W_{1},\Sigma_{1})$ is 2-dimensional and so $\mathrm{vcd}(W_{1})=\mathrm{cd}_{\mathbb{Q}}(W_{1})=2$. Moreover, for $n\geq 2$, as $W_{n}\cong W_{1}\ast\tilde{A}_{n}$ the Mayer–Vietoris sequence (cf. [20, Theorem 2]) yields $\mathrm{cd}_{\mathbb{Q}}(W_{n})=\max\{\mathrm{cd}_{\mathbb{Q}}(W_{1}),\mathrm{cd}_{\mathbb{Q}}(\tilde{A}_{n})\}=n$. Therefore, $\mathrm{cd}_{\mathbb{Q}}(\bar{G}_{1})=2$ and, for all $n\geq 2$, $\mathrm{cd}_{\mathbb{Q}}(\bar{G}_{n})=n$.
From Proposition 4.4 the following question arises.
Question 1.
For which Coxeter groups does $\mathrm{alg}\text{-}\mathrm{rk}(W)=\mathrm{cd}_{\mathbb{Q}}(W)$ hold?
For instance, every affine Coxeter group $(W,S)$ has algebraic rank $|S|-1$ and so the equality $\mathrm{alg}\text{-}\mathrm{rk}(W)=\mathrm{cd}_{\mathbb{Q}}(W)$ holds. The product $W_{1}\times W_{2}\times\cdots\times W_{n}=:\Gamma$ of finitely many Coxeter groups satisfies $\mathrm{alg}\text{-}\mathrm{rk}(\Gamma)=\mathrm{cd}_{\mathbb{Q}}(\Gamma)$ whenever $\mathrm{alg}\text{-}\mathrm{rk}(W_{i})=\mathrm{cd}_{\mathbb{Q}}(W_{i})$ for every $i=1,\ldots,n$ (cf. [5, Theorem 5.6]). Another example is provided by M. W. Davis and T. Januszkiewicz in [24]: for each right-angled Artin group $A(\Gamma)$ there is a right-angled Coxeter group $W_{\Gamma}$ which contains $A(\Gamma)$ as a subgroup of finite index. In particular, $\mathrm{cd}_{\mathbb{Q}}(W_{\Gamma})=\mathrm{cd}_{\mathbb{Q}}(A(\Gamma))$ and $\mathrm{alg}\text{-}\mathrm{rk}(W_{\Gamma})=\mathrm{alg}\text{-}\mathrm{rk}(A(\Gamma))$. It suffices now to notice that $\mathrm{alg}\text{-}\mathrm{rk}(A(\Gamma))$ coincides with the largest size of a clique of $\Gamma$, which is also the dimension of the associated Salvetti complex; see for example [34, Theorems 3.4 and 3.5]. In particular, $\mathrm{alg}\text{-}\mathrm{rk}(W_{\Gamma})=\mathrm{cd}_{\mathbb{Q}}(W_{\Gamma})=\mathrm{vcd}(W_{\Gamma})$. It is well-known that there are Coxeter groups satisfying $\mathrm{vcd}(W)\neq\mathrm{cd}_{\mathbb{Q}}(W)$ (cf. [4, Remarks (3)]). Therefore, apart from Question 1, also the following question arises.
Question 2.
For which Coxeter groups $(W,S)$ does $\mathrm{cd}_{\mathbb{Q}}(W)=\mathrm{vcd}(W)$ hold?
A class of examples answering Question 2 affirmatively is given by Coxeter groups of virtual cohomological dimension $\leq 2$. In the following, we answer Question 1 and Question 2 for hyperbolic Coxeter groups.
4.1. Hyperbolic Coxeter groups
Let $(W,S)$ be a Coxeter group and put $|S|=n$. Given $V=\mathrm{span}_{\mathbb{R}}\{\alpha_{s}\mid s\in S\}$, let $B\colon V\times V\to\mathbb{R}$ be the symmetric bilinear form on $V$ given by
$B(\alpha_{s},\alpha_{t})=-\cos\Big{(}\frac{\pi}{m_{st}}\Big{)},\quad\,\forall\,s,t\in S.$
Let $\sigma\colon W\to\mathrm{GL}(V)$ denote the geometric representation of $W$ (cf. [31, §5.3]). The bilinear form $B$ is invariant under $\mathrm{im}(\sigma)$ (cf. [31, Proposition 5.3]). Define the cone of $(V,B)$ by
(4.5) $C:=\{\lambda\in V\mid B(\lambda,\alpha_{s})<0\ \text{for all $s\in S$}\}.$
The irreducible Coxeter group $(W,S)$ is said to be hyperbolic if
  • (a)
    $B$ has signature $(n-1,1)$, and
  • (b)
    $B(\lambda,\lambda)<0$ for all $\lambda\in C$.
Hyperbolic Coxeter groups exist only for $3\leq n\leq 10$ (cf. [31, p. 141]). For a hyperbolic Coxeter group $(W,S)$ the image of the geometric representation is contained in $O(n-1,1)$. In particular, $\mathrm{im}(\sigma)$ is a lattice in $O(n-1,1)$ if, and only if, $(W,S)$ is hyperbolic (cf. [31, Remark 6.8], [8, pp. 131-135] and [35]). If $(W,S)$ is hyperbolic and crystallographic as a Coxeter group, then $\mathrm{im}(\sigma)$ also stabilises some lattice $\Lambda\subset V$ (cf. [31, §6.6]). Thus, if =$\mathscr{B}=\{\lambda_{i}\mid 1\leq i\leq|S|\}$ is a free generating system of $\Lambda$, there exists some positive integer $e\in\mathbb{N}$ such that $B(\lambda_{i},\lambda_{j})\in\tfrac{1}{e}\mathbb{Z}$. Defining $\mathbf{G}(\mathbb{Z})=\{\alpha\in\mathrm{GL}_{\mathbb{Z}}(\Lambda)\mid B(\alpha(v_{i}),\alpha(v_{j}))=B(v_{i},v_{j})\}$ yields an affine algebraic group scheme $\mathbf{G}$ defined over $\mathbb{Z}$ satisfying
(4.6) $W\subseteq\mathbf{G}(\mathbb{Z})\subset\mathbf{G}(\mathbb{R})=O(n-1,1).$
Since $W$ is a lattice in $O(n-1,1)$, so is $\mathbf{G}(\mathbb{Z})$ and $|\mathbf{G}(\mathbb{Z})\colon W|<\infty$. Consequently,
Fact 4.6.
Let $(W,S)$ be a hyperbolic and crystallographic Coxeter group. Then $W$ is an arithmetic lattice in $O(n-1,1)$, where $n=|S|$.
One refers to $(W,S)$ as compact hyperbolic Coxeter group if $W$ is cocompact in $O(n-1,1)$ (cf. [31, §6.9]).
Proposition 4.7.
Let $(W,S)$ be a compact hyperbolic Coxeter group. Then
$\mathrm{alg}\text{-}\mathrm{rk}(W)=1\quad\text{and}\quad\mathrm{cd}_{\mathbb{Q}}(W)=\mathrm{vcd}(W)=|S|-1\geq 2.$
Proof.
By [31, §6.8], every proper special subgroup of $(W,S)$ is spherical. This fact has two consequences. First, from [42, Theorem 17.1(iii)$\Rightarrow$(ii)] one deduces that $\mathrm{alg}\text{-}\mathrm{rk}(W)=1$. Second, the poset of all spherical special subgroups of $(W,S)$ is isomorphic to the poset of all spherical special subgroups of an affine Coxeter group $(W^{\prime},S^{\prime})$ with $|S^{\prime}|=|S|$, where the two posets are ordered by inclusion. Therefore, $B_{R}(W,S)$ and $B_{R}(W^{\prime},S^{\prime})$ (for $R\in\{\mathbb{Z},\mathbb{Q}\}$) have the same dimension. By [4], one concludes that
$\mathrm{cd}_{\mathbb{Q}}(W)=\mathrm{cd}_{\mathbb{Q}}(W^{\prime})=|S|-1=\mathrm{vcd}(W^{\prime})=\mathrm{vcd}(W).$
Remark 4.8.
The statement of Proposition 4.7 can be alternatively deduced from [7, Corollary 11.4.3] and the fact that every cocompact discrete subgroup of $O(n-1,1)$ is Gromov-hyperbolic (cf. [29, p. 93]).
Example 4.1.
From Lannér’s classification (cf. [31, §6.9]), [31, Proposition 6.6] and [31, §6.7], the Coxeter group associated to the Coxeter diagram
[…]
is the unique compact hyperbolic crystallographic Coxeter group of rank greater than or equal to $4$. Therefore, by Proposition 4.7, if $G$ is a group with a locally finite twin root datum of type $(W,S)$, the geometric completion $\bar{G}$ of $G$ satisfies:
$31=\mathrm{alg}\text{-}\mathrm{rk}(W)=\mathrm{flat}\text{-}\mathrm{rk}(\bar{G})\quad\text{and}\quad\mathrm{cd}_{\mathbb{Q}}(\bar{G})=\mathrm{cd}_{\mathbb{Q}}(W)=\mathrm{vcd}(W)=$ 3 .
Proposition 4.9.
Let $(W,S)$ be a non-compact hyperbolic Coxeter group. Then
$\mathrm{vcd}(W)=\mathrm{alg}\text{-}\mathrm{rk}(W)=|S|-2.$
In particular, $\mathrm{cd}_{\mathbb{Q}}(W)=|S|-2$.
Proof.
Since $\mathrm{alg}\text{-}\mathrm{rk}(W)\leq\mathrm{cd}_{\mathbb{Q}}(W)\leq\mathrm{vcd}(W)$ (cf. Theorem 4.2 and Section 2.3), it suffices to prove the first part of the statement. Throughout the proof, put $|S|=n$. By [31, p. 140], every non-compact hyperbolic Coxeter group $(W,S)$ contains an affine special subgroup $(W^{\prime},S^{\prime})$ with $|S^{\prime}|=n-1$. Hence,
(4.7) $\mathrm{vcd}(W)\geq\mathrm{alg}\text{-}\mathrm{rk}(W)\geq\mathrm{alg}\text{-}\mathrm{rk}(W^{\prime})=n-2.$
It remains to show that $\mathrm{vcd}(W)\leq n-2$. We first deal with the crystallographic case. Let $X$ denote the symmetric space associated to $O(n-1,1)$. Then,
(4.8) $\dim(X)=\binom{n}{2}-\binom{n-1}{2}=n-1.$
Since $W\backslash X$ is non-compact, one has $\partial X\not=\emptyset$ (cf. [7, Theorem 9.3]). Equivalently, the $\mathbb{Q}$-rank of $\mathbf{G}$ satisfies $\ell(\mathbf{G})\geq 1$, which yields
$\mathrm{vcd}(W)=\dim(X)-\ell(\mathbf{G})\leq n-2,$
(cf. [7, Corollary 11.4.3]).
In the remaining non-crystallographic cases (cf. [31, §6.9, Figure 3]), Fact 4.10 below applies and the Bestvina’s complex $B_{\mathbb{Z}}(W,S)$ has at most $n-1$ vertices. By [4], one concludes that $\mathrm{vcd}(W)\leq n-2$. ∎
Fact 4.10.
Let $(W,S)$ be a non-compact hyperbolic Coxeter group. Suppose that either $|S|=3$ or $(W,S)$ is non-crystallographic with $|S|\geq 4$. Then $(W,S)$ admits exactly $|S|-1$ maximal spherical subsets.
Proof.
Let first $|S|=3$. Since $(W,S)$ is irreducible, its Coxeter diagram is a segment of length $2$ with edges $\{a,b\}$ and $\{b,c\}$. Since $(W,S)$ is a non-compact hyperbolic Coxeter group, at least one of the two edges, say $\{a,b\}$, is labelled with $\infty$ (cf. [31, p. 140]). If $m_{bc}=\infty$, the maximal spherical subsets of $(W,S)$ are $\{a,c\}$ and $\{b\}$. Otherwise, they are $\{a,c\}$ and $\{b,c\}$.
In what follows, by an $r$-subset of $S$ we mean a subset with $r$ elements. Moreover, we make use of the classifications of spherical, affine and hyperbolic Coxeter groups by their Coxeter diagram as given, for example, in [31, pp. 32,34,142-144]). Since $(W,S)$ is a hyperbolic Coxeter group, observe that every $r$-subset of $S$, with $r\leq|S|-2$, is spherical (cf. [31, Proposition 6.8], recalling that every proper special subgroup of a Coxeter group of positive type is spherical). Thus the maximal spherical subsets of $S$ have cardinality either $|S|-1$ or $|S|-2$.
Let now $(W,S)$ be non-crystallographic with $|S|\geq 4$. By [31, §6.9, Figure 3], there are only seven cases: six with $|S|=4$ and one with $|S|=6$. If $S=\{a,b,c,d\}$, then $(W,S)$ has one of following Coxeter diagram:
[…]
Let first $(W,S)$ be associated to one of the Coxeter diagrams in (4.9). Among the $3$-subsets of $S$, only $\{a,c,d\}$ and $\{b,c,d\}$ are spherical. Note also that $\{a,b\}$ is the only $2$-subset of $S$ which is not contained in both $\{a,c,d\}$ and $\{b,c,d\}$. Thus, $\{a,c,d\}$ and $\{b,c,d\}$ and $\{a,b\}$ are the maximal spherical subsets of $S$. Assume now that $(W,S)$ is associated to one of the Coxeter diagrams in (4.10). Among the $3$-subsets of $S$, the following ones are spherical: $\{a,b,d\}$, $\{a,c,d\}$ and $\{b,c,d\}$. Hence, they are the maximal spherical subsets of $S$.
Finally, let $S=\{a,b,c,d,e,f\}$. Then $(W,S)$ has the following Coxeter diagram:
[…]
For $x\in S$, the set $S\setminus\{x\}$ is spherical if and only if $x\not\in\{a,b\}$. In particular, the $4$-subset $S\setminus\{x,y\}$ is contained in a spherical $5$-subset of $S$ if, and only if, either $x\not\in\{a,b\}$ or $y\not\in\{a,b\}$. One concludes that the maximal spherical subsets of $S$ are $S\setminus\{x\}$, for some $x\not\in\{a,b\}$, and $S\setminus\{a,b\}$. ∎
Remark 4.11.
Fact 4.10 is not true for arbitrary non-compact hyperbolic crystallographic Coxeter groups. For example, let $(W,S)$ be the Coxeter group defined by the following Coxeter diagram:
[…]
Among the $3$-subsets of $S=\{a,b,c,d\}$, only $\{a,c,d\}$ is spherical. Moreover, there are three spherical $2$-subsets of $S$, namely $\{a,b\}$, $\{b,c\}$ and $\{b,d\}$, that are not contained in $\{a,c,d\}$. As a conclusion, $(W,S)$ admits exactly $4=|S|$ maximal spherical subsets and so does not satisfy the conclusion of Fact 4.10.
4.2. Coxeter groups with maximal algebraic rank
For every Coxeter group $(W,S)$, one has
(4.11) $\mathrm{alg}\text{-}\mathrm{rk}(W)\leq\mathrm{cd}_{\mathbb{Q}}(W)\leq\mathrm{vcd}(W)\leq|S|-1.$
The first inequality comes from the fact that the rank of a free abelian group is its rational cohomological dimension, and from the fact that $H\leq G$ implies $\mathrm{cd}_{\mathbb{Q}}(H)\leq\mathrm{cd}_{\mathbb{Q}}(G)$. For the second and the third inequalities, see Section 2.3.
The maximum possible algebraic rank for $(W,S)$, namely $|S|-1$, is attained only in the affine case as shown below.
Proposition 4.12.
Let $(W,S)$ be a Coxeter group with $|S|\geq 2$. Then $(W,S)$ is affine if, and only if, $\mathrm{alg}\text{-}\mathrm{rk}(W)=|S|-1$.
Proof.
Assume that $(W,S)$ is affine. Then $W$ is the semidirect product of its translation subgroup $T$ by a spherical Coxeter group (cf. [31, Proposition 4.2]). In particular, $T$ is a free abelian subgroup of rank $|S|-1$ and $\mathrm{alg}\text{-}\mathrm{rk}(W)=|S|-1$ (cf. [41, Proposition 2.3]).
Now assume that $\mathrm{alg}\text{-}\mathrm{rk}(W)=|S|-1$. Since $|S|-1>0$, $(W,S)$ is non-spherical. By Krammer’s theorem (cf. [37, Theorem 6.8.3]), the algebraic rank of $W$ coincides with the rank of a so-called standard free abelian subgroup, say $\prod_{j=1}^{k}H_{j}$ (cf. [37, p. 154]). In particular, there are irreducible non-spherical subsets $I_{1},\ldots,I_{k}$ of $S$ which are pairwise perpendicular (i.e., for all $j\neq l$, every element of $I_{j}$ commutes with every element of $I_{l}$) and such that
(4.12) $|S|-1=\mathrm{alg}\text{-}\mathrm{rk}(W)=\sum_{j=1}^{k}\mathrm{alg}\text{-}\mathrm{rk}(H_{j})\leq\sum_{j=1}^{k}(|I_{j}|-1).$
As the $I_{j}$’s are pairwise disjoint, $|I_{1}|+\ldots+|I_{k}|\leq|S|$ and, by (4.12), one has
$|S|-1\leq\sum_{j=1}^{k}|I_{j}|-k\leq|S|-k.$
Necessarily, $k=1$ and $I_{1}=S$. Since $|S|\geq 2$, the algebraic rank of $H_{1}$ is greater than $1$ and then $H_{1}$ is the translation subgroup of the affine special subgroup $(W_{I_{1}},I_{1})$ (cf. [37, p 154]). One concludes that $(W,S)=(W_{I_{1}},I_{1})$ is affine. ∎
Corollary 4.13.
Let $(W,S)$ be an affine Coxeter group. Then $\mathrm{alg}\text{-}\mathrm{rk}(W)=\mathrm{vcd}(W)=|S|-1$.
4.3. Coxeter groups of rank $3$
Given a Coxeter group $(W,S)$ of rank $3$, we completely describe how the virtual cohomological dimension and the rational cohomological dimension of $W$ relate to its algebraic rank.
Proposition 4.14.
Let $(W,S)$ be a Coxeter group with $|S|=3$, say $S=\{a,b,c\}$, and with Coxeter matrix $[m_{st}]_{s,t\in S}$. Then $(W,S)$ falls into one of the following cases:
  • (i)
    $\mathrm{alg}\text{-}\mathrm{rk}(W)=\mathrm{cd}_{\mathbb{Q}}(W)=\mathrm{vcd}(W)=0$. It is attained if, and only if, $(W,S)$ is spherical. I.e., $(W,S)$ has of one of the following types: $A_{1}^{3}$, $A_{3}$, $B_{3}$, $H_{3}$ or $I_{2}(m)\times A_{1}$ for $3\leq m<\infty$);
  • (ii)
    $\mathrm{alg}\text{-}\mathrm{rk}(W)=\mathrm{cd}_{\mathbb{Q}}(W)=\mathrm{vcd}(W)=2$. It is attained if, and only if, $(W,S)$ is affine. I.e., $(W,S)$ has of one of the following types: $\widetilde{A}_{2}$, $\widetilde{B}_{2}=\widetilde{C}_{2}$ or $\widetilde{G}_{2}$;
  • (iii)
    $\mathrm{alg}\text{-}\mathrm{rk}(W)=\mathrm{cd}_{\mathbb{Q}}(W)=\mathrm{vcd}(W)=1$. It is attained if, and only if, at least one among $m_{ab}$, $m_{bc}$ and $m_{ca}$ is $\infty$;
  • (iv)
    $\mathrm{alg}\text{-}\mathrm{rk}(W)=1<\mathrm{cd}_{\mathbb{Q}}(W)=\mathrm{vcd}(W)=2$. I.e., $(W,S)$ is compact hyperbolic of rank 3.
Proof.
By (4.11), one has $\mathrm{alg}\text{-}\mathrm{rk}(W)\leq\mathrm{cd}_{\mathbb{Q}}(W)\leq\mathrm{vcd}(W)\leq 2$. The Coxeter group $(W,S)$ is spherical if, and only if, $\mathrm{alg}\text{-}\mathrm{rk}(W)=0$ or, equivalently, $\mathrm{vcd}(W)=0$. By Proposition 4.12, $\mathrm{alg}\text{-}\mathrm{rk}(W)=2$ if, and only if, $(W,S)$ is affine. Therefore, Corollary 4.13 yields (ii). By [31, §6.7], (iv) is a consequence of Proposition 4.7 and items (i),(ii) and (iii).
Finally, assume $\mathrm{alg}\text{-}\mathrm{rk}(W)=1$. By Stallings–Swan theorem, both the conditions $\mathrm{cd}_{\mathbb{Q}}(W)=1$ and $\mathrm{vcd}(W)=1$ are equivalent to $W$ being infinite and virtually free. In turn, by [40, Theorem 34], $W$ is infinite and virtually free if, and only if, $W$ is infinite, $\Gamma_{\infty}(W,S)$ is chordal and every clique in $\Gamma_{\infty}(W,S)$ is spherical. Since $|S|=3$, this is equivalent to require that at least one among $m_{ab}$, $m_{bc}$ and $m_{ca}$ is $\infty$. ∎
Example 4.2.
The Coxeter groups associated to the Coxeter diagrams
[…]
are the only compact hyperbolic and crystallographic Coxeter groups of rank $3$. Let $(W,S)$ be one as above. If $G$ is a group with a locally finite twin root datum of type $(W,S)$, the geometric completion $\bar{G}$ of $G$ satisfies $\mathrm{flat}\text{-}\mathrm{rk}(\bar{G})=1$ and $\mathrm{cd}_{\mathbb{Q}}(\bar{G})=2.$
5. $\infty$-decompositions and Haglund–Paulin trees
Let $\Delta=(\mathcal{C},\delta)$ be a building of type $(W,S)$. For $T\subseteq S$ and $\mathcal{M}\subseteq\mathcal{C}$, set
(5.1) $\mathcal{R}_{T}(\mathcal{M}):=\{\mathrm{Res}_{T}(c)\mid c\in\mathcal{M}\},$
(cf. Section 2.4). If necessary, one writes $\mathrm{Res}_{T}^{\Delta}(c)$ and $\mathcal{R}_{T}^{\Delta}(\mathcal{M})$ in place of $\mathrm{Res}_{T}(c)$ and $\mathcal{R}_{T}(\mathcal{M})$, respectively.
Definition 5.1.
Let $(W,S)$ be a Coxeter group.
  1. (a)
    A graph of special subgroups $(\mathcal{G},\Lambda)$ of $(W,S)$ based on a connected graph $\Lambda$ is a graph of groups assigning to each $\lambda\in\mathscr{V}(\Lambda)\sqcup\mathscr{E}(\Lambda)$ the special subgroup $\mathcal{G}_{\lambda}=W_{S_{\lambda}}$ of $(W,S)$, and for every $e\in\mathscr{E}(\Lambda)$ the monomorphism $\mathcal{G}_{e}\hookrightarrow\mathcal{G}_{t(e)}$ is the inclusion map. In case that the group homomorphism $\tau\colon\pi_{1}(\mathcal{G},\Lambda)\to W$ induced by the inclusion maps $W_{S_{\lambda}}\hookrightarrow W$, $\lambda\in\mathscr{V}(\Lambda)\sqcup\mathscr{E}(\Lambda)$, is an isomorphism, the pair $(\mathcal{G},\Lambda)$ is said to be a visual graph of groups decomposition of $(W,S)$ (cf. [40]).
  2. (b)
    Let $\Delta=(\mathcal{C},\delta)$ be a building of type $(W,S)$ and $(\mathcal{G},\Lambda)$ be a graph of special subgroups of $(W,S)$ as in (a). For every non-empty subset $\mathcal{M}\subseteq\mathcal{C}$ define a graph $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{M})=(\mathscr{V}(\mathcal{M}),\mathscr{E}(\mathcal{M}))$ in the sense of J-P. Serre as follows:
    $\mathscr{V}(\mathcal{M}):=\bigsqcup_{v\in\mathscr{V}(\Lambda)}\mathcal{R}_{S_{v}}(\mathcal{M})\quad\text{and}\quad\mathscr{E}(\mathcal{M}):=\bigsqcup_{e\in\mathscr{E}(\Lambda)}\mathcal{R}_{S_{e}}(\mathcal{M}).$
    For every $e\in\mathscr{E}(\Lambda)$, although $S_{e}=S_{\bar{e}}$, here $\mathcal{R}_{S_{e}}(\mathcal{M})$ and $\mathcal{R}_{S_{\bar{e}}}(\mathcal{M})$ are regarded as disjoint copies of the same set. The edge-reversing, the origin and the terminus maps are defined, respectively, as
    (5.2) $\overline{\mathrm{Res}_{S_{e}}(c)}:=\mathrm{Res}_{S_{\bar{e}}}(c),\quad o(\mathrm{Res}_{S_{e}}(c)):=\mathrm{Res}_{S_{o(e)}}(c),\quad t(\mathrm{Res}_{S_{e}}(c)):=\mathrm{Res}_{S_{t(e)}}(c),$
    for all $e\in\mathscr{E}(\Lambda)$ and $c\in\mathcal{M}$.
Remark 5.1.
Since $W$ is generated by involutions, if $(\mathcal{G},\Lambda)$ is a visual graph of groups decomposition, then $\Lambda$ must be a tree.
Remark 5.2.
  • (i)
    The assignments in (5.2) are well-defined. For the inversion map it is obvious. Moreover, for $e\in\mathscr{E}(\Lambda)$ and $v\in\{o(e),t(e)\}$, observe that $\mathrm{Res}_{S_{e}}(c)=\mathrm{Res}_{S_{e}}(d)$ implies $\delta(c,d)\in W_{S_{e}}\subseteq W_{S_{v}}$ and then, by (2.5), $\mathrm{Res}_{S_{v}}(c)=\mathrm{Res}_{S_{v}}(d)$.
  • (ii)
    If $\Lambda$ is a combinatorial graph, so is $\Gamma=\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{M})$. In other terms, if the map $e\in\mathscr{E}(\Lambda)\longmapsto(o(e),t(e))\in\mathscr{V}(\Lambda)\times\mathscr{V}(\Lambda)$ is injective, so is the map $\mathrm{Res}_{S_{e}}(c)\in\mathscr{E}(\Gamma)\longmapsto(\mathrm{Res}_{S_{o(e)}}(c),\mathrm{Res}_{S_{t(e)}}(c))\in\mathscr{V}(\Gamma)\times\mathscr{V}(\Gamma)$. Indeed, $\mathrm{Res}_{S_{o(e)}}(c)\cap\mathrm{Res}_{S_{t(e)}}(c)=\mathrm{Res}_{S_{e}}(c)$ for all $e\in\mathscr{E}(\Lambda)$ and $c\in\mathcal{M}$.
Example 5.1.
[…]
  • (i)
    The graph $\Gamma_{(\mathcal{G},\Lambda)}^{\Sigma}(W)$ is isomorphic to the graph $X=X(W;W_{S_{\blacktriangledown}},W_{S_{\blacktriangle}})$ defined in [49, §I.4, proof of Theorem 7]. By [49, §I.4, Theorems 6 and 7], $\Gamma_{(\mathcal{G},\Lambda)}^{\Sigma}(W)$ is a tree if, and only if, $W\simeq W_{S_{\blacktriangledown}}\amalg_{W_{S_{\bullet}}}W_{S_{\blacktriangle}}$.
  • (ii)
    More generally, let $\Delta=(\mathcal{C},\delta)$ be a building of type $(W,S)$. Then $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C})$ coincides with the graph introduced by F. Haglund and F. Paulin in [30, §4.2]. By [30, p. 146, Remarques] and [30, Lemme 4.3] (or Corollary 5.8 below), $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C})$ is a tree if, and only if, $W\simeq W_{S_{\blacktriangledown}}\amalg_{W_{S_{\bullet}}}W_{S_{\blacktriangle}}$. From now on, the graph $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C})$ constructed here will be denoted by $\Gamma_{\blacktriangledown\!/\!\blacktriangle}^{\Delta}(\mathcal{C})$.
Lemma 5.3.
Let $\Delta=(\mathcal{C},\delta)$ be a building of type $(W,S)$, let $\mathcal{M}\subseteq\mathcal{C}$ be non-empty and $(\mathcal{G},\Lambda)$ be a graph of special subgroups of $(W,S)$. Denote by $\Sigma(W,S)$ the standard Coxeter complex of type $(W,S)$.
  • (i)
    Every $W$-isometry $\alpha\colon W\longrightarrow\mathcal{C}$ induces a graph isomorphism
    $\varphi_{\alpha}:\Gamma_{(\mathcal{G},\Lambda)}^{\Sigma(W,S)}(W)\longrightarrow\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\alpha(W))$
    mapping $wW_{S_{x}}$ to $\mathrm{Res}_{S_{x}}^{\Delta}(\alpha(w))$, for all $x\in\mathscr{V}(\Lambda)\sqcup\mathscr{E}(\Lambda)$ and $w\in W$.
  • (ii)
    For $c\in\mathcal{C}$, let $\mathfrak{A}_{c}$ be the set of all apartments in an atlas of $\Delta$ (cf. Fact 2.1) containing $c$ as a chamber. Then,
    $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C})=\bigcup_{\Sigma\in\mathfrak{A}_{c}}\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C}(\Sigma)).$
Proof.
Part (ii) is straightforward from $\mathcal{C}=\bigcup_{\Sigma\in\mathfrak{A}_{c}}\mathcal{C}(\Sigma)$ (cf. Fact 2.1(ii)). Thus it suffices to prove (i). For every $x\in\mathscr{V}(\Lambda)\sqcup\mathscr{E}(\Lambda)$, the assignment
$\varphi_{\alpha,x}:wW_{S_{x}}\longmapsto\mathrm{Res}_{S_{x}}^{\Delta}(\alpha(w)),\quad\forall\,w\in W$
yields a well-defined bijection from $\mathcal{R}_{S_{x}}^{\Sigma(W,S)}(W)$ to $\mathcal{R}_{S_{x}}^{\Delta}(\alpha(W))$. Indeed, $\varphi_{\alpha,x}$ is well-defined and injective by the following:
$\begin{array}[]{ccl}w_{1}W_{S_{x}}=w_{2}W_{S_{x}}&\Longleftrightarrow&\delta_{W}(w_{1},w_{2})=\delta(\alpha(w_{1}),\alpha(w_{2}))\in W_{S_{x}}\\ &\Longleftrightarrow&\mathrm{Res}_{S_{x}}^{\Delta}(\alpha(w_{1}))=\mathrm{Res}_{S_{x}}^{\Delta}(\alpha(w_{2})),\end{array}$
for all $w_{1},w_{2}\in W$ (cf. (2.5)). Finally, the surjectivity of $\varphi_{\alpha,x}$ is part of the definition. ∎
Lemma 5.4.
Let $\Delta=(\mathcal{C},\delta)$ be a building of type $(W,S)$, $\mathcal{M}\subseteq\mathcal{C}$ be non-empty, and $(\mathcal{G},\Lambda)$ be a graph of special subgroups of $(W,S)$. Assume that $S=\bigcup_{v\in\mathscr{V}(\Lambda)}S_{v}$. Let $\gamma=(c_{0},c_{1},\ldots,c_{n})$ be a gallery in $\Delta$ with $c_{i}\in\mathcal{M}$, for every $0\leq i\leq n$. Then, for all $v,w\in\mathscr{V}(\Lambda)$, $\gamma$ induces a path $\mathfrak{p}_{\gamma,v\rightarrow w}$ in $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{M})$ from $\mathrm{Res}_{S_{v}}(c_{0})$ to $\mathrm{Res}_{S_{w}}(c_{n})$.
Proof.
It suffices to show the statement for $\mathscr{E}(\Lambda)\neq\emptyset$ and for a gallery $\gamma=(c,d)$ of length $2$. That is, given $c,d\in\mathcal{M}$ with $\delta(c,d)=s\in S$ and $v,w\in\mathscr{V}(\Lambda)$, there is a path in $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{M})$ from $\mathrm{Res}_{S_{v}}(c)$ to $\mathrm{Res}_{S_{w}}(d)$. By hypothesis, there is $z\in\mathscr{V}(\Lambda)$ such that $s\in S_{z}$, and $\mathrm{Res}_{S_{z}}(c)=\mathrm{Res}_{S_{z}}(d)$ (cf. (2.5)). If $v=w=z$, the path $\mathfrak{p}_{\gamma,v\to w}$ is the trivial path. In all the remaining cases, let $(e_{1},\ldots,e_{n})$ be a path in $\Lambda$ from $v$ to $w$ running through $z$, i.e., $o(e_{m})=z$ for some $1\leq m\leq n$. If $m=1$, then $v=z$ and $\mathrm{Res}_{S_{v}}(c)=\mathrm{Res}_{S_{z}}(c)=\mathrm{Res}_{S_{z}}(d)$ is the origin of $\mathrm{Res}_{S_{e_{1}}}(d)$ in $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{M})$ (cf. (5.2)). Thus the sequence
$(\mathrm{Res}_{S_{e_{1}}}(d),\ldots,\mathrm{Res}_{S_{e_{n}}}(d))$
gives a path in $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{M})$ from $\mathrm{Res}_{S_{v}}(c)$ to $\mathrm{Res}_{S_{w}}(d)$. Similarly, if $m\geq 2$ one has $t(e_{m-1})=z=o(e_{m})$ and then
$\mathrm{Res}_{S_{t(e_{m-1})}}(c)=\mathrm{Res}_{S_{z}}(c)=\mathrm{Res}_{S_{z}}(d)=\mathrm{Res}_{S_{o(e_{m})}}(d).$
Hence, the sequence
$(\mathrm{Res}_{S_{e_{1}}}(c),\ldots,\mathrm{Res}_{S_{e_{m-1}}}(c),\mathrm{Res}_{S_{e_{m}}}(d),\ldots,\mathrm{Res}_{S_{e_{n}}}(d))$
defines a path in $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{M})$ from $\mathrm{Res}_{S_{v}}(c)$ to $\mathrm{Res}_{S_{w}}(d)$. ∎
Let $\Delta=(\mathcal{C},\delta)$ be a building of type $(W,S)$. A non-empty subset $\mathcal{M}\subseteq\mathcal{C}$ is gallery connected if, for all $c,d\in\mathcal{M}$, there exists a gallery $(c_{0}=c,c_{1},\dots,c_{k}=d)$ in $\Delta$ with $c_{i}\in\mathcal{M}$ for all $0\leq i\leq k$ (cf. [1, Definition 5.43]). For example, the set $\mathcal{C}$ is gallery connected (cf. [1, Example 5.44(a)]).
Lemma 5.4 implies the following.
Corollary 5.5.
Let $\Delta=(\mathcal{C},\delta)$ be a building of type $(W,S)$ and $(\mathcal{G},\Lambda)$ be a graph of special subgroups of $(W,S)$. Assume that $S=\bigcup_{v\in\mathscr{V}(\Lambda)}S_{v}$. Then, for every gallery connected subset $\mathcal{M}\subseteq\mathcal{C}$, the graph $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{M})$ is connected. In particular, $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C})$ is connected.
Moreover, let $\{\Sigma_{i}=(\mathcal{C}(\Sigma_{i}),\delta|_{\Sigma_{i}})\}_{i\in I}$ be a collection of apartments of $\Delta$ with $\bigcap_{i\in I}\mathcal{C}(\Sigma_{i})\neq\emptyset$. Then,
$\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}\Bigg{(}\bigcap_{i\in I}\mathcal{C}(\Sigma_{i})\Bigg{)}=\bigcap_{i\in I}\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C}(\Sigma_{i})),$
$\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}\Bigg{(}\bigcup_{i\in I}\mathcal{C}(\Sigma_{i})\Bigg{)}=\bigcup_{i\in I}\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C}(\Sigma_{i})),$
and $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}\Big{(}\bigcap_{i\in I}\mathcal{C}(\Sigma_{i})\Big{)}$ and $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}\Big{(}\bigcup_{i\in I}\mathcal{C}(\Sigma_{i})\Big{)}$ are both connected.
Proof.
For the first part of the statement, apply Lemma 5.4. The second part of the statement follows from the first part and the fact that, since $\bigcap_{i\in I}\mathcal{C}(\Sigma_{i})\neq\emptyset$, the sets $\bigcap_{i\in I}\mathcal{C}(\Sigma_{i})$ and $\bigcup_{i\in I}\mathcal{C}(\Sigma_{i})$ are gallery connected (cf. [1, Example 5.44(c)]). ∎
Proposition 5.6.
Let $\Delta=(\mathcal{C},\delta)$ be a building of type $(W,S)$ and $(\mathcal{G},\Lambda)$ be a graph of special subgroups of $(W,S)$. Assume that $S=\bigcup_{v\in\mathscr{V}(\Lambda)}S_{v}$. Then $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C})$ is a tree if, and only if, $\Gamma_{(\mathcal{G},\Lambda)}^{\Sigma(W,S)}(W)$ is a tree.
Proof.
Let first $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C})$ be a tree and $\alpha:W\longrightarrow\mathcal{C}$ be a $W$-isometry. Hence, $\alpha(W)$ defines an apartment of $\Delta$. By Corollary 5.5 one has that $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\alpha(W))$ is connected and thus a tree. By Lemma 5.3(i), also $\Gamma_{(\mathcal{G},\Lambda)}^{\Sigma(W,S)}(W)$ is a tree.
Suppose conversely that $\Gamma_{(\mathcal{G},\Lambda)}^{\Sigma(W,S)}(W)$ is a tree. For $c\in\mathcal{C}$, let $\mathfrak{A}_{c}$ be the collection of all the apartments in some atlas of $\Delta$ having $c$ as a chamber. By Lemma 5.3, for every $\Sigma\in\mathfrak{A}_{c}$ the graph $\Gamma_{\Sigma}:=\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C}(\Sigma))$ is a subtree of $\Gamma:=\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C})$, and $\Gamma=\bigcup_{\Sigma\in\mathfrak{A}_{c}}\Gamma_{\Sigma}$. In order to conclude that $\Gamma$ is a tree, it is sufficient to show that, for every finite non-empty subset $\mathcal{S}\subseteq\mathfrak{A}_{c}$, the graph $\Gamma_{\mathcal{S}}:=\bigcup_{\Sigma\in\mathcal{S}}\Gamma_{\Sigma}$ is a subtree of $\Gamma$. Indeed, every closed path in $\Gamma$ without backtrackings must be contained in the union of finitely many $\Gamma_{\Sigma}$’s. One proceeds by induction on the cardinality of $\mathcal{S}$. For $|\mathcal{S}|=1$ there is nothing to prove. Let now $\mathcal{S}\subseteq\mathfrak{A}_{c}$ with $|\mathcal{S}|\geq 2$ and suppose that the claim holds for all subsets of $\mathfrak{A}_{c}$ of cardinality $|\mathcal{S}|-1$. Let $\Sigma\in\mathcal{S}$ and $\mathcal{S}^{\prime}:=\mathcal{S}\setminus\{\Sigma\}$. By induction, both $\Gamma_{\Sigma}$ and $\Gamma_{\mathcal{S}^{\prime}}$ are trees. Moreover, by Corollary 5.5,
$\Gamma_{\Sigma}\cap\Gamma_{\mathcal{S}^{\prime}}=\bigcup_{\Sigma^{\prime}\in\mathcal{S}}(\Gamma_{\Sigma}\cap\Gamma_{\Sigma^{\prime}})$
is a non-empty connected subgraph of $\Gamma_{\Sigma}$ and thus is a tree. By van Kampen’s theorem, $\Gamma_{\mathcal{S}}=\Gamma_{\Sigma}\cup\Gamma_{\mathcal{S}^{\prime}}$ is a tree. ∎
The following theorem generalises results of F. Haglund and F. Paulin (cf. Corollary 5.8) to arbitrary visual graph of groups decompositions.
Theorem 5.7.
Let $\Delta=(\mathcal{C},\delta)$ be a building of type $(W,S)$ and $(\mathcal{G},\Lambda)$ be a graph of special subgroups of $(W,S)$. Then the following conditions are equivalent:
  • (i)
    $(\mathcal{G},\Lambda)$ is a visual graph of groups decomposition of $(W,S)$;
  • (ii)
    $S=\bigcup_{v\in\mathscr{V}(\Lambda)}S_{v}$ and $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C})$ is a tree.
Proof.
Note that if $(\mathcal{G},\Lambda)$ is a visual graph of groups decomposition of $(W,S)$, then $W$ is generated by the union of the vertex groups and therefore $S=\bigcup_{v\in\mathscr{V}(\Lambda)}S_{v}$. The statement is a consequence of Proposition 5.6 and the main theorem of Bass–Serre theory (cf. [49, §I.5.4, Theorem 13]), because (i) is equivalent to the fact that $\Gamma_{(\mathcal{G},\Lambda)}^{\Sigma(W,S)}(W)$ is a tree. Indeed, as $\mathcal{R}_{T}^{\Sigma(W,S)}(W)=W/W_{T}$ for $T\subseteq S$, the graph $\Gamma_{(\mathcal{G},\Lambda)}^{\Sigma(W,S)}(W)$ is the universal Bass–Serre graph of $(\mathcal{G},\Lambda)$ (cf. [49, p. 51]). ∎
Corollary 5.8 ([30, Lemme 4.3, Remarques at p. 146]).
Let $\Delta=(\mathcal{C},\delta)$ be a building of type $(W,S)$ and $S_{\blacktriangledown},S_{\blacktriangle}\subseteq S$ such that $S=S_{\blacktriangledown}\cup S_{\blacktriangle}$. Put $S_{\bullet}=S_{\blacktriangledown}\cap S_{\blacktriangle}$ and let $\Gamma_{\blacktriangledown\!/\!\blacktriangle}^{\Delta}(\mathcal{C})$ be as in Example 5.1(ii). Then the following conditions are equivalent:
  • (i)
    $S=S_{\blacktriangledown}\cup S_{\blacktriangle}$ is an $\infty$-decomposition of $(W,S)$;
  • (ii)
    $W\simeq W_{S_{\blacktriangledown}}\amalg_{W_{S_{\bullet}}}W_{S_{\blacktriangle}}$;
  • (iii)
    $\Gamma_{\blacktriangledown\!/\!\blacktriangle}^{\Delta}(\mathcal{C})$ is a tree.
Proof.
The statement follows by combining Theorem 5.7 and the fact that $S=S_{\blacktriangledown}\cup S_{\blacktriangle}$ is an $\infty$-decomposition of $(W,S)$ if, and only if, $W\simeq W_{S_{\blacktriangledown}}\amalg_{W_{S_{\bullet}}}W_{S_{\blacktriangle}}$ (cf. Section 1). ∎
Let $\Delta=(\mathcal{G},\delta)$ be a building of type $(W,S)$ and $(\mathcal{G},\Lambda)$ be a graph of special subgroups of $(W,S)$. Then every group $G$ acting on $\Delta$ has an induced action on $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{M})$, for every (non-empty) $G$-invariant subset $\mathcal{M}\subseteq\mathcal{C}$.
Corollary 5.9.
Let $G$ be a $\sigma$-compact t.d.l.c. group acting transitively on a building $\Delta=(\mathcal{C},\delta)$ of type $(W,S)$, and let $c\in\mathcal{C}$. Suppose that $(W,S)$ admits a visual graph of groups decomposition $(\mathcal{G},\Lambda)$. Then $G$ is topologically isomorphic to the fundamental group of the tree of t.d.l.c. groups $(\widetilde{\mathcal{G}},\Lambda)$, where $\widetilde{\mathcal{G}}_{\lambda}$ is the setwise stabiliser of $\mathrm{Res}_{S_{\lambda}}(c)$ for every $\lambda\in\mathscr{V}(\Lambda)\sqcup\mathscr{E}(\Lambda)$, and the continuous open monomorphisms $\widetilde{\mathcal{G}}_{e}\hookrightarrow\widetilde{\mathcal{G}}_{t(e)}$, $e\in\mathscr{E}(\Lambda)$, are given by inclusion. Moreover, the universal Bass–Serre tree of $\pi_{1}(\widetilde{\mathcal{G}},\Lambda)$ is isomorphic to $\Gamma_{(\widetilde{\mathcal{G}},\Lambda)}^{\Delta}(\mathcal{C})$.
Proof.
By Theorem 5.7, $\Gamma=\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C})$ is a tree. Since $G$ acts chamber-transitively on $\Delta$, one has that $g\cdot\mathrm{Res}_{T}(c)=\mathrm{Res}_{T}(g\cdot c)$ for all $g\in G$ and $T\subseteq S$. In particular, for every $\lambda\in\mathscr{V}(\Lambda)\sqcup\mathscr{E}(\Lambda)$, the group $G$ acts transitively and with open stabilisers on $\mathcal{R}_{S_{\lambda}}(\mathcal{C})$. Therefore, $G$ acts on $\Gamma$ with open stabilisers and $G\backslash\Gamma=\Lambda$. Let $(\widetilde{\mathcal{G}},\Lambda)$ be the graph of t.d.l.c. groups defined in the statement. By the main theorem of Bass–Serre theory, the homomorphism $\varphi\colon\pi_{1}(\widetilde{\mathcal{G}},\Lambda)\to G$ induced by inclusion maps $\widetilde{\mathcal{G}}_{\lambda}\hookrightarrow G$, $\lambda\in\mathscr{V}(\Lambda)\sqcup\mathscr{E}(\Lambda)$, is an isomorphism of abstract groups. Moreover, the universal Bass–Serre tree of $\pi_{1}(\widetilde{\mathcal{G}},\Lambda)$ is isomorphic to $\Gamma_{(\widetilde{\mathcal{G}},\Lambda)}^{\Delta}(\mathcal{C})$. It remains to prove that $\varphi$ is open and has an open inverse. Recall that the group topology of $\pi_{1}(\widetilde{\mathcal{G}},\Lambda)$ has a neighbourhood basis at $1$ formed by the open subgroups of the vertex-groups of $(\widetilde{\mathcal{G}},\Lambda)$. In particular, $\pi_{1}(\widetilde{\mathcal{G}},\Lambda)$ is a t.d.l.c. group and $\varphi$ is open. Moreover, by the open mapping theorem (cf. [53, Theorem 6.19]), the homomorphism $\varphi^{-1}$ is open as well. ∎
6. Spherical $\infty$-decompositions and the number of ends
6.1. $\infty$-Decompositions for a given subset $J$
Given a Coxeter group $(W,S)$ and a subset $J\subsetneq S$, it is possible to investigate $\infty$-decompositions $S=S_{\blacktriangledown}\cup S_{\blacktriangle}$ satisfying $J=S_{\blacktriangledown}\cap S_{\blacktriangle}$ considering the presentation diagram $\Gamma_{\infty}(W,S)$. Let $\Xi_{J}$ be the subgraph of $\Gamma_{\infty}(W,S)$ spanned by $S\setminus J$, i.e.,
(6.1) $\mathscr{V}(\Xi_{J})$ $=S\setminus J,$
$\mathscr{E}(\Xi_{J})$ =$\{\{u,v\}\in\mathscr{P}_{2}(S\setminus J)\mid m_{uv}\not=\infty\},$
where $\mathscr{P}_{2}(S\setminus J)$ is the set of all the subsets of $S\setminus J$ of cardinality $2$. Then there exists a non-trivial $\infty$-decomposition $S=S_{\blacktriangledown}\cup S_{\blacktriangle}$ for $(W,S)$ satisfying $S_{\blacktriangledown}\cap S_{\blacktriangledown}=J$ if, and only if, $\Xi_{J}$ is disconnected. Indeed, let $L\subseteq S\setminus J$ be the set of vertices of a connected component of $\Xi_{J}$ and let $V=S\setminus(J\cup L)$. Then $S_{\blacktriangledown}=L\cup J$ and $S_{\blacktriangle}=V\cup J$ define a non-trivial $\infty$-decomposition of $(W,S)$. Consequently, one has the following.
Fact 6.1.
Let $(W,S)$ be a Coxeter group. For every $J\subsetneq S$, the following are equivalent:
  • (i)
    there exists a non-trivial $\infty$-decomposition $S=S_{\blacktriangledown}\cup S_{\blacktriangle}$ of $(W,S)$ satisfying $J=S_{\blacktriangledown}\cap S_{\blacktriangle}$;
  • (ii)
    $\Xi_{J}$ has at least $2$ connected components.
6.2. Coxeter groups with more than one end
We now state a result due to M. W. Davis in terms of spherical $\infty$-decompositions of a Coxeter group $(W,S)$, i.e., $\infty$-decompositions $S=S_{\blacktriangledown}\cup S_{\blacktriangle}$ of $(W,S)$ in which $S_{\blacktriangledown}\cap S_{\blacktriangle}$ is a spherical subset of $(W,S)$. A similar reformulation has also been provided in [40, Corollary 16] but with a different approach. Recall that an $\infty$-decomposition $S=S_{\blacktriangledown}\cup S_{\blacktriangle}$ of $(W,S)$ is said to be non-trivial if $S_{\blacktriangledown}\cap S_{\blacktriangle}$ is a proper subset of $S_{\blacktriangledown}$ and $S_{\blacktriangle}$.
Theorem 6.2 ([22, Theorem 8.7.2] and [40, Corollary 16]).
Let $(W,S)$ be a Coxeter group. Then the following are equivalent:
  • (i)
    the group $W$ has more than one end;
  • (ii)
    $(W,S)$ has a non-trivial spherical $\infty$-decomposition.
Proof.
The implication (ii)$\Rightarrow$(i) follows by Stallings’ decomposition theorem. Indeed, assume that $(W,S)$ has a non-trivial spherical $\infty$-decomposition $S=S_{\blacktriangledown}\cup S_{\blacktriangle}$. Then, it is evident from the Coxeter presentation that $W$ is a non-trivial free product of the special subgroups corresponding to $S_{\blacktriangledown}$ and $S_{\blacktriangle}$ amalgamated over the special subgroup generated by $S_{\blacktriangledown}\cap S_{\blacktriangle}$ (cf. [30, 40]).
Conversely, suppose that the Coxeter group $(W,S)$ has more than one end. Then one must have $\mathrm{H}^{1}(W,\mathbb{Q}W)\not=0$ (cf. [26, Chapter IV, Theorem 6.10]). By [22, Corollary 8.5.2],
(6.2) $\mathrm{H}^{1}(W,\mathbb{Q}W)\simeq\bigoplus_{J\subset S,\atop{J\text{ spherical}}}\mathbb{Q}W^{J}\otimes\mathrm{H}^{1}(K,K^{S\setminus J};\mathbb{Q}),$
where $K$ is the fundamental chamber of $|\Sigma(W,S)_{\mathrm{Dav}}$| (cf. [22, p.  126]) and $\mathbb{Q}W^{J}$ is the $\mathbb{Q}$-vector space spanned by the set $W^{J}$ (cf. [22, Definition 4.7.4]). In particular, there exists $J\subseteq S$ spherical such that $\mathrm{H}^{1}(K,K^{S\setminus J};\mathbb{Q})\not=0$. By Künneth’s theorem, $\mathrm{H}_{1}(K,K^{S\setminus J};\mathbb{Q})\not=0$. As $K$ is contractible (cf. [22, Lemma  7.2.5(i)]), one has a short exact sequence
(6.3) $0\longrightarrow\mathrm{H}_{1}(K,K^{S\setminus J};\mathbb{Q})\longrightarrow\mathrm{H}_{0}(K^{S\setminus J},\mathbb{Q})\longrightarrow\mathbb{Q}\longrightarrow 0$
Thus, $\mathrm{H}_{1}(K,K^{S\setminus J};\mathbb{Q})\not=0$ implies that $K^{S\setminus J}$ has more than one path-connected component. By definition, $K^{S\setminus J}$ is the union of the connected spaces $K_{s}$, $s\in S\setminus J$, where each $K_{s}$ is the geometric realisation of the poset $\{T\subseteq S\mid T\text{ spherical, }s\in T\}$ (cf. [22, §7.2]). Moreover, $K_{s}\cap K_{t}\not=\emptyset$ if, and only if, $m_{st}<\infty$. Hence, since $K^{S\setminus J}$ has at least two connected components, the number of connected components of $\Xi_{J}$ is at least two. Then Fact 6.1 yields the claim. ∎
Let $e(W)$ denote the number of ends of the Coxeter group $W$. Since $W$ is finitely generated, one has $e(W)\in\{0,1,2,\infty\}$. Moreover, $e(W)=0$ if and only if $W$ is finite. The following is a direct consequence of Theorem 6.2.
Corollary 6.3.
Let $(W,S)$ be a Coxeter group with Coxeter matrix $[m_{st}]_{s,t\in S}$.
  • (i)
    Let $(W,S)$ be right-angled (i.e., $m_{st}\in\{2,\infty\}$ for all $s\neq t$). Then $e(W)\geq 2$ if, and only if, there exists a non-trivial $\infty$-decomposition $S=S_{\blacktriangledown}\cup S_{\blacktriangle}$ such that $m_{st}\leq 2$ for all $s,t\in S_{\blacktriangledown}\cap S_{\blacktriangle}$.
  • (ii)
    If $m_{st}<\infty$ for all $s,t\in S$, then $e(W)\leq 1$. In particular,
    • (iia)
      if $(W,S)$ is a hyperbolic Coxeter group with $|S|\geq 4$ (cf. [31, §6.9]), then $e(W)=1$;
    • (iib)
      if $(W,S)$ is affine (cf. [31, §6.5]) then $e(W)\in\{1,2\}$. Moreover, $e(W)=2$ if and only if $W\simeq D_{\infty}=C_{2}\amalg C_{2}$.
6.3. Completely spherical $\infty$-decompositions
By a repeated application of Theorem 6.2, M. W. Davis observed that every Coxeter group $(W,S)$ is isomorphic to the fundamental group $\pi_{1}(\mathcal{G},\mathscr{T})$ of a finite tree of groups, in which each vertex-group is a special subgroup of $W$ with at most one end, and each edge-group is a spherical special subgroup (cf. [22, Proposition 8.8.2]). The injection of each edge-group into its endpoint vertex-group is given by inclusion, and the isomorphism between the fundamental group and $W$ is induced by the inclusion maps of the vertex-groups into $W$. In other words, every Coxeter group $(W,S)$ admits a visual graph decomposition $(\mathcal{G},\mathscr{T})$ in the sense of [40] with spherical edge-groups and vertex-groups with at most one end. If $W$ is infinite and every vertex-group of $(\mathcal{G},\mathscr{T})$ is a spherical special subgroup of $W$, we say that $(\mathcal{G},\mathscr{T})$ is a completely spherical $\infty$-decomposition of $(W,S)$. If $(\mathcal{H},\mathscr{A})$ is another completely spherical $\infty$-decomposition of $(W,S)$, then there exists a bijection $\phi$ between the vertices of $\mathscr{T}$ and the vertices of $\mathscr{A}$ such that, for each vertex $v$ of $\mathscr{T}$, the vertex-group $\mathcal{G}_{v}$ is conjugate to the vertex-group $\mathcal{H}_{\phi(v)}$ (cf. [40, Proof of Theorem 19]).
Combining Theorem 6.2, the main results of [27], and [40, Theorem 34], one obtains the following version of the Karrass–Pietrowski–Solitar theorem for Coxeter groups.
Theorem 6.4.
Let $(W,S)$ be a Coxeter group. Then the following is equivalent:
  • (i)
    $W$ is infinite and virtually free;
  • (ii)
    $\mathrm{vcd}(W)=1$ (or, equivalently, $\mathrm{cd}_{\mathbb{Q}}(W)=1$);
  • (iii)
    $(W,S)$ admits a completely spherical $\infty$-decomposition;
  • (iv)
    $(W,S)$ is not spherical, but every clique $\Lambda\subseteq\Gamma_{\infty}(W,S)$ is spherical, and $\Gamma_{\infty}(W,S)$ is chordal.
Recall that an undirected graph $\Gamma=(V,E)$ is said to be chordal if the only induced cycles in $\Gamma$ have $3$ vertices. An induced cycle is a vertex-sequence $\gamma=(v_{0},v_{1},\ldots,v_{n},v_{0})$ with $n\geq 2$ and such that $\{v_{i},v_{j}\}\in E$ if, and only if, $v_{i}$ and $v_{j}$ are adjacent terms in the sequence $\gamma$.
Corollary 6.5.
Let $G$ be a t.d.l.c. group acting chamber-transitively on a locally finite building $\Delta$ with compact stabilisers. If $\mathrm{cd}_{\mathbb{Q}}(G)=1$ then $G$ decomposes as fundamental group of a finite tree of profinite groups and $G$ is unimodular.
Proof.
By Theorems 3.4 and 6.4, $\mathrm{cd}_{\mathbb{Q}}(W)=1$ and $W\simeq\pi_{1}(\mathcal{G},\Lambda)$ for some finite tree of spherical special subgroups $(\mathcal{G},\Lambda)$ of $(W,S)$. By Corollary 5.9 and since $G$ acts with compact stabilisers on $\Delta$, $G$ acts on $\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C})$ with compact open vertex stabilisers and $G\backslash\Gamma_{(\mathcal{G},\Lambda)}^{\Delta}(\mathcal{C})=\Lambda$. The fact that $G$ is unimodular is due to [2, Propositions 1.2 and 3.6]. ∎
6.4. Completely-$\mathscr{P}$infty-decompositions
Let $\Gamma=(V,E)$ be a finite undirected graph. The graph of maximal cliques $\mathbf{mx}(\Gamma)=(V(\mathbf{mx}(\Gamma)),E(\mathbf{mx}(\Gamma)))$ is given by
(6.4) $V(\mathbf{mx}(\Gamma))$ =$=\{\Lambda\subseteq\Gamma\mid\Lambda\ \text{a maximal clique}\}$
$E(\mathbf{mx}(\Gamma))$ =$\big{\{}\{\Lambda,\Xi\}\subseteq\mathscr{P}_{2}(V(\mathbf{mx}(\Gamma)))\mid\ V(\Lambda)\cap V(\Xi)\not=\emptyset\big{\}}.$
Here $\mathscr{P}_{2}(V(\mathbf{mx}(\Gamma)))$ denotes the set of all subsets of $V(\mathbf{mx}(\Gamma))$ of size $2$. Let $\overrightarrow{\mathbf{mx}}(\Gamma)$ be the associated combinatorial graph in the sense of J-P. Serre (cf. §2.1). A maximal subtree $\mathscr{T}\subseteq\overrightarrow{\mathbf{mx}}(\Gamma)$ has the clique intersection property if, for every pair $(\Lambda,\Xi)$ of maximal cliques in $\Gamma$, the unique path without backtrackings $(\Lambda=\Xi_{0},\Xi_{1},\ldots,\Xi_{r}=\Xi)$ in $\mathscr{T}$ satisfies $\Lambda\cap\Xi\subseteq\Xi_{i}$ for all $0\leq i\leq r$. The existence of a maximal subtree $\mathscr{T}\subseteq\overrightarrow{\mathbf{mx}}(\Gamma)$ satisfying this property is equivalent to require that $\Gamma$ is chordal (cf. [6, Proposition 7.2]).
For a Coxeter group $(W,S)$ one says that $\Gamma_{\infty}(W,S)$ has the $\mathscr{P}$-clique property if, for every clique $(S^{\prime},\mathscr{P}_{2}(S^{\prime}))$ of $\Gamma_{\infty}(W,S)$, the Coxeter group $(W_{S^{\prime}},S^{\prime})$ has the property $\mathscr{P}$.
Theorem 6.6.
Let $(W,S)$ be an infinite Coxeter group such that $\Gamma_{\infty}(W,S)$ has the $\mathscr{P}$-clique property. If $\Gamma_{\infty}(W,S)$ is chordal, then $W$ admits a completely-$\mathscr{P}$infty-decomposition $\pi_{1}(\mathcal{G},\mathscr{T})$ (i.e., $W\simeq\pi_{1}(\mathcal{G},\mathscr{T})$ where $(\mathcal{G},\mathscr{T})$ is a tree of special subgroups such that all vertex-groups and edge-groups satisfy property $\mathscr{P}$).
In particular, $\mathscr{T}$ can be chosen to be a maximal subtree of $\overrightarrow{\mathbf{mx}}(\Gamma_{\infty}(W,S))$ satisfying the clique intersection property.
Proof.
By [6, Proposition 7.2] there exists a maximal subtree $\mathscr{T}\subseteq\overrightarrow{\mathbf{mx}}(\Gamma_{\infty}(W,S))$ with the clique intersection property. Build the graph of special subgroups $(\mathcal{G},\mathscr{T})$ as follows: each vertex-group $\mathcal{G}_{\Lambda}$, $\Lambda\in V(\mathscr{T})$, is the special subgroup $W_{V(\Lambda)}$ of $(W,S)$ generated by the set of vertices $V(\Lambda)$ of the maximal clique $\Lambda$, each edge-group $\mathcal{G}_{e}=\mathcal{G}_{\bar{e}}$, with $e=\{\Lambda,\Xi\}\in E(\mathscr{T})$, is the intersection $\mathcal{G}_{o(e)}\cap\mathcal{G}_{t(e)}$, and the monomorphisms are given by inclusions $\iota_{e}\colon\mathcal{G}_{e}\hookrightarrow\mathcal{G}_{o(e)}$, for every $e\in E(\mathscr{T})$. One must prove that $W\simeq\pi_{1}(\mathcal{G},\mathscr{T})$. To this end, for $J\subseteq S$ denote by $\langle J\mid\mathcal{R}_{J}\rangle$ the Coxeter presentation of $(W_{J},J)$. Thus $\pi_{1}(\mathcal{G},\mathscr{T})$ admits the presentation:
$\Bigg{\langle}\bigsqcup_{\Lambda\in V(\mathscr{T})}V(\Lambda)\,\Bigg{|}\,\bigsqcup_{\Lambda\in V(\mathscr{T})}\mathcal{R}_{V(\Lambda)},v_{\Lambda}=v_{\Xi}\ \forall v\in V(\Lambda)\cap V(\Xi)\ \forall\{\Lambda,\Xi\}\in E(\mathscr{T})\Bigg{\rangle},$
where $v_{\Lambda}$ and $v_{\Xi}$ denote the copies of the vertex $v\in V(\Lambda)\cap V(\Xi)$ in the sets $V(\Lambda)$ and $V(\Xi)$, respectively. Since each $s\in S$ is a vertex of some maximal clique of $\Gamma_{\infty}(W,S)$, one has that $S=\bigcup_{\Lambda\in V(\mathscr{T})}V(\Lambda)$. Similarly, $\mathcal{R}_{S}=\bigcup_{\Lambda\in V(\mathscr{T})}\mathcal{R}_{V(\Lambda)}$ because each edge of $\Gamma_{\infty}(W,S)$ belongs to a maximal clique. Therefore, it is sufficient to prove that the disjoint unions appearing in the above presentation can be relaxed to unions (i.e., $\pi_{1}(\mathcal{G},\mathscr{T})$ admits the presentation $\langle S\mid\mathcal{R}_{S}\rangle$ and so it is isomorphic to $W$). To this end, for every $\{\Lambda,\Xi\}\in E({\mathbf{mx}}(\Gamma_{\infty}(W,S)))$ denote by $\mathfrak{p}_{\{\Lambda,\Xi\}}$ the unique path without backtracking in $\mathscr{T}$ between $\Lambda$ and $\Xi$, and note that the images $\iota_{e}(V(\Lambda)\cap V(\Xi))$ and $\iota_{f}(V(\Lambda)\cap V(\Xi))$ are pointwise identified for all $e,f\in\mathfrak{p}_{\{\Lambda,\Xi\}}$ since $\mathscr{T}$ has the clique intersection property. ∎
Being finite can be considered as a property $\mathscr{P}$. In this case, one can use Theorem 6.6 to give an alternative proof of the implication (iv)$\Rightarrow$(iii) in Theorem 6.4. Another example of a suitable property $\mathscr{P}$ is the following: a group $G$ is said to be slender (or Noetherian) if every subgroup of $G$ is finitely generated. By Theorem 6.6, one deduces the following result due to O. Varghese (cf. [54, Corollary 3.5]).
Corollary 6.7.
Let $(W,S)$ be a Coxeter group. If $\Gamma_{\infty}(W,S)$ is chordal and has the slender-clique property, then $(W,S)$ is coherent (i.e., every finitely generated subgroup is finitely presented).
Proof.
By Theorem 6.6, $(W,S)$ has a completely-slender $\infty$-decomposition. By [54, Corollary 3.5], every slender Coxeter group is coherent. The claim follows by the fact that the free product of coherent groups amalgamated over a slender subgroup is coherent (cf. [33, Theorem 8]). ∎
7. The number of ends of a compactly generated t.d.l.c. group
7.1. Cayley–Abels graphs
Given a t.d.l.c. group $G$, a Cayley–Abels graph of $G$ is defined to be a connected locally finite graph $\Gamma$ on which $G$ acts vertex-transitively with compact open stabilisers. It is well-known that $G$ admits a Cayley–Abels graph if, and only if, $G$ is compactly generated. Moreover, all Cayley–Abels graphs of a compactly generated t.d.l.c. group $G$ are quasi-isometric to each other (cf. [16, 38, 55] for instance).
From now on, $G$ will denote a compactly generated t.d.l.c. group.
Example 7.1.
Given a building $\Delta=(\mathcal{C},\delta)$ of type $(W,S)$, the chamber graph $(\mathcal{C},E(\mathcal{C}))$ of $\Delta$ (cf. [36, Definition 4.1]) is the undirected graph whose set of vertices is $\mathcal{C}$, and $\{c,d\}\in E(\mathcal{C})$ if, and only if, $\delta(c,d)\in S$. Assume that $\Delta$ is locally finite and $G$ acts transitively on $\Delta$ with compact open stabilisers. Then $(\mathcal{C},E(\mathcal{C}))$ is a Cayley–Abels graph of $G$.
Following [19], a compact generating system of $G$ is a pair $(K,S)$ such that
  • $K$ is a compact open subgroup of $G$;
  • $S$ is a finite subset of $G$ satisfying $S=S^{-1}$ and $S\cap K=\emptyset$;
  • $G$ is algebraically generated by $K\cup S$.
Given a compact generating system $(K,S)$ of $G$, the associated Cayley–Abels graph $\Gamma(G,S,K)$ is the connected graph in the sense of J-P. Serre consisting of the following data:
$\mathscr{V}(\Gamma(G,S,K))=G/K\quad\text{and}\quad\mathscr{E}(\Gamma(G,S,K))=\{(gK,gsK)\mid g\in G,s\in S\}.$
The origin map is given by the projection on the first coordinate, the terminus map is given by the projection on the second coordinate, while the edge inversion map permutes the first and second coordinates.
7.2. Almost $(G,K)$-invariant sets
Let $K$ be a compact open subgroup of $G$. Following [15, 27], a set $B\subseteq G/K$ is said to be almost $(G,K)$-invariant if, for all $g\in G$, $gB=_{a}B$ (i.e., the symmetric difference is finite) and, for all $k\in K$, $kB=B$. Note that a finite set $B\subseteq G/K$ is almost $(G,K)$-invariant if, and only if, $B=B_{1}\sqcup\ldots\sqcup B_{n}$ where each $B_{i}$ denotes the set of all left $K$-cosets that are necessary to cover a single $K$-double-coset lying in the preimage of $B$ in $G$.
Example 7.2.
Let $G$ be a compactly generated t.d.l.c. group, let $(K,S)$ be a compact generating system and $\Gamma=\Gamma(G,K,S)$ the associated Cayley–Abels graph. Given a finite connected subgraph $\Gamma^{\prime}$ of $\Gamma$, denote by $\Gamma-\Gamma^{\prime}$ the subgraph spanned by all the vertices of $\Gamma$ that are not in $\Gamma^{\prime}$. Every connected component $C$ of $\Gamma-\Gamma^{\prime}$ determines the almost $(G,K)$-invariant set
(7.1) $B_{C}:=\{gK\in G/K\mid g^{-1}K\in\mathscr{V}(C)\}.$
Clearly, $kB_{C}=B_{C}$ for every $k\in K$. Therefore, since $G$ is algebraically generated by $K\cup S$, it suffices to prove that $sB_{C}=_{a}B_{C}$ for every $s\in S$. Given $s\in S$, assume that $gK\in B_{C}$ and $sgK\notin B_{C}$. It means that $g^{-1}s^{-1}K$ belongs to the boundary of $C$, i.e., $g^{-1}s^{-1}K$ is not a vertex of $C$ but there is an edge in $\Gamma$ connecting $g^{-1}s^{-1}K$ to a vertex of $C$ (which in this case is $g^{-1}K$). As $\Gamma^{\prime}$ is finite, the boundary of $C$ is finite. Since $S^{-1}=S$, one concludes that $sB_{C}=_{a}B_{C}$.
7.3. Almost $(G,K)$-invariant functions
Let $(K,S)$ be a compact generating system of $G$. In this section we recall the relation between the rational discrete first-degree cohomology of $G$ and the almost $(G,K)$-invariant sets. The $\mathbb{Q}$-vector space $\mathrm{Hom}_{\mathbb{Q}}(\mathbb{Q}[G/K],\mathbb{Q})$ admits the left $G$-action given by
$g\cdot\alpha(x)=\alpha(g^{-1}x),\quad\alpha\in\mathrm{Hom}_{\mathbb{Q}}(\mathbb{Q}[G/K],\mathbb{Q}),g\in G,x\in G/K.$
One writes $\alpha=_{a}\beta$ if $\alpha(x)=\beta(x)$ for all but finitely many elements $x\in G/K$. An element $\alpha\in\mathrm{Hom}_{\mathbb{Q}}(\mathbb{Q}[G/K],\mathbb{Q})$ is said to be almost $(G,K)$-invariant if, for all $g\in G$, $g\cdot\alpha=_{a}\alpha$ and, for all $k\in K$, $k\cdot\alpha=\alpha$. For example, the characteristic function $\chi_{B}$ of an almost $(G,K)$-invariant set $B$ is an almost $(G,K)$-invariant function. By [15, Proposition 3.15], for every compact open subgroup $K$ of a t.d.l.c. group $G$ one has
(7.2) $\mathrm{dH}^{1}(G,\mathbb{Q}[G/K])\cong\frac{\mathrm{AInv}_{K}(G,\mathbb{Q})}{C(G/K)+\mathbb{Q}[G/K]^{K}},$
[…]
A locally finite connected graph $\Gamma$ has at least $n$ ends if there is a finite subgraph $\Gamma^{\prime}$ such that $\Gamma-\Gamma^{\prime}$ has at least $n$ infinite connected components. The graph $\Gamma$ has $n$ ends if it has at least $n$ ends but not at least $n+1$ ends. Moreover, $\Gamma$ has infinitely many ends if, for all $n\in\mathbb{N}$, $\Gamma$ has at least $n$ ends. The number of ends $e(G)$ of $G$ is defined to be the number of ends of a(ny) Cayley–Abels graph of $G$. Note that this number is well-defined since the Cayley–Abels graphs of $G$ are quasi-isometric to each other. It is well-known that $e(G)\in\{0,1,2,\infty\}$ and that $e(G)=0$ if, and only if, $G$ is compact.
Theorem 7.1.
Let $G$ be a non-compact compactly generated t.d.l.c. group and let $K$ be a compact open subgroup. The $\mathbb{Q}$-vector space $\mathrm{dH}^{1}(G,\mathbb{Q}[G/K])$ has dimension $e(G)-1$ if $G$ has finitely many ends. Moreover, $G$ has infinitely many ends if and only if $\mathrm{dH}^{1}(G,\mathbb{Q}[G/K])$ is infinite-dimensional.
Proof.
The argument is based on the proof of [50, Satz IV]. Let $(K,S)$ be a compact generating system of $G$. It suffices to prove the following statements, for all $n,m\geq 1$: if the associated Cayley–Abels graph $\Gamma(G,K,S)$ has at least $n$ ends, then the dimension of $\mathrm{dH}^{1}(G,\mathbb{Q}[G/K])$ is at least $n-1$; if the dimension of $\mathrm{dH}^{1}(G,\mathbb{Q}[G/K])$ is at least $m$, then $\Gamma(G,K,S)$ has at least $m+1$ ends.
Assume that $\Gamma(G,K,S)$ has at least $n$ ends. Then there exists a finite subgraph $\Gamma^{\prime}$ such that $\Gamma(G,K,S)-\Gamma^{\prime}$ has at least $n$ infinite connected components $C_{1},\ldots,C_{n}$. Let $B_{C_{1}},\ldots,B_{C_{n}}$ be the infinite almost $(G,K)$-invariant subsets determined by $C_{1},\ldots,C_{n}$, respectively (cf. Example 7.2). Observe now that the characteristic functions $\chi_{{}_{B_{C_{1}}}},\ldots,\chi_{{}_{B_{C_{n-1}}}}$ are linearly independent modulo $C(G/K)+\mathbb{Q}[G/K]^{K}$. Indeed, any $\mathbb{Q}$-linear combination $t_{1}\chi_{{}_{B_{C_{1}}}}+\cdots+t_{n-1}\chi_{{}_{B_{C_{n-1}}}}$ vanishes (at least) on the infinite set $B_{C_{n}}$. Since $B_{C_{i}}\cap B_{C_{j}}=\emptyset$ for every $i\neq j$, the only $\mathbb{Q}$-linear combination $t_{1}\chi_{{}_{B_{C_{1}}}}+\cdots+t_{n-1}\chi_{{}_{B_{C_{n-1}}}}$ that is constant almost everywhere is the one with $t_{1}=\ldots=t_{n-1}=0$. By (7.2), one concludes that the dimension of $\mathrm{dH}^{1}(G,\mathbb{Q}[G/K])$ is at least $n-1$.
For every $B\subseteq G/K$ let $C_{B}=\{gK\in G/K\mid g^{-1}K\in B\}$. One checks that $B$ is finite if, and only if, $C_{B}$ is finite. Assume that the dimension of $\mathrm{dH}^{1}(G,\mathbb{Q}[G/K])$ is at least $m$. By (7.2) and [15, Lemma 4.4], there exist $m$ almost $(G,K)$-invariant subsets $B_{1},\dots,B_{m}$ of $G/K$ whose characteristic functions $\chi_{B_{1}},\ldots,\chi_{B_{m}}$ are linearly independent modulo $C(G/K)+\mathbb{Q}[G/K]^{K}$. Denote by $\delta C_{B_{i}}$ the set of vertices in $\Gamma(G,K,S)$ which are not in $C_{B_{i}}$ but are adjacent to a vertex of $C_{B_{i}}$. As $B_{i}$ is infinite and almost $(G,K)$-invariant, $C_{B_{i}}$ is infinite and $\delta C_{B_{i}}$ is finite (for the finiteness of $\delta C_{B_{i}}$ one can argue as in Example 7.2). Let $\Gamma^{\prime}$ be the subgraph of $\Gamma(G,K,S)$ spanned by $\delta C_{B_{1}}\cup\ldots\cup\delta C_{B_{m}}$, and denote by $D_{1},\ldots,D_{p}$ the sets of vertices of the $p$ infinite connected components of $\Gamma(G,K,S)-\Gamma^{\prime}$, respectively. Observe that
$G/K=\mathscr{V}(\Gamma(G,K,S))=F\sqcup\bigsqcup_{j=1}^{p}D_{j},$
where $F$ is the (finite) set collecting all the elements of $\delta C_{B_{1}}\cup\ldots\cup\delta C_{B_{m}}$ and all vertices of any finite connected component of $\Gamma(G,K,S)-\Gamma^{\prime}$. In particular, for every $1\leq i\leq m$ one has
(7.3) $C_{B_{i}}=(C_{B_{i}}\cap F)\sqcup\bigsqcup_{j=1}^{p}(C_{B_{i}}\cap D_{j}).$
We claim that, for all $1\leq i\leq m$ and $1\leq j\leq p$, either $C_{B_{i}}\supseteq D_{j}$ or $C_{B_{i}}\cap D_{j}=\emptyset$. Suppose indeed that there is $gK\in D_{j}\setminus C_{B_{i}}$. Since $D_{j}\cap\delta C_{B_{i}}=\emptyset$, every adjacent vertex of any $hK\in D_{j}$ does not belong to $C_{B_{i}}$. Hence, every path in $\Gamma(G,K,S)$ with vertices in $D_{j}$ and starting at $gK$ has all its vertices in $D_{j}\setminus C_{B_{i}}$. Since $D_{j}$ spans a connected subgraph of $\Gamma(G,K,S)$, every vertex in $D_{j}$ can be connected to $gK$ by such a path. Then $D_{j}=D_{j}\setminus C_{B_{i}}$, which yields $D_{j}\cap C_{B_{i}}=\emptyset$. Therefore, by (7.3), for every $1\leq i\leq m$ there is $\{j_{1},\ldots,j_{r_{i}}\}\subseteq\{1,\ldots,p\}$ such that
(7.4) $\chi_{C_{B_{i}}}=\chi_{C_{B_{i}}\cap F}+\sum_{k=1}^{r_{i}}\chi_{D_{j_{k}}}.$
Denote by $\mathcal{A}\mathcal{C}(G/K,\mathbb{Q})$ the space of all functions $f\colon G/K\longrightarrow\mathbb{Q}$ which are constant almost everywhere, i.e., $f$ is constant on all but finitely many elements of $G/K$. Moreover, $\mathcal{A}\mathcal{C}(G/K,\mathbb{Q})$ is regarded as a subspace of the space $\mathcal{F}(G/K,\mathbb{Q})$ of all maps from $G/K$ to $\mathbb{Q}$. Since $\chi_{C_{B_{i}}\cap F}\in\mathcal{A}\mathcal{C}(G/K,\mathbb{Q})$ for every $1\leq i\leq m$, (7.4) implies that the $\mathbb{Q}$-subspace $\mathcal{S}_{1}$ in $\mathcal{F}(G/K,\mathbb{Q})/\mathcal{A}\mathcal{C}(G/K,\mathbb{Q})$ generated by $\{\chi_{C_{B_{i}}}+\mathcal{A}\mathcal{C}(G/K,\mathbb{Q})\}_{i=1}^{m}$ is contained in the $\mathbb{Q}$-subspace $\mathcal{S}_{2}$ generated by $\{\chi_{D_{j}}+\mathcal{A}\mathcal{C}(G/K,\mathbb{Q})\}_{j=1}^{p}$. Note that $\dim_{\mathbb{Q}}\mathcal{S}_{1}=m$. Indeed, $\chi_{B_{i}},\ldots,\chi_{B_{m}}$ are linearly independent modulo $C(G/K)+\mathbb{Q}[G/K]^{K}$ (and so modulo $\mathcal{A}\mathcal{C}(G/K,\mathbb{Q})$) and, for all $t_{1},\ldots,t_{m}\in\mathbb{Q}$ and $x\in G$, one has
$\Big{(}t_{1}\chi_{C_{B_{1}}}+\ldots+t_{m}\chi_{C_{B_{m}}}\Big{)}(xK)=\Big{(}t_{1}\chi_{B_{1}}+\ldots+t_{m}\chi_{B_{m}}\Big{)}(x^{-1}K).$
Moreover, arguing as in the first part of the proof, $\chi_{D_{1}},\ldots,\chi_{D_{p-1}}$ are linearly independent modulo $\mathcal{A}\mathcal{C}(G/K,\mathbb{Q})$. However, $\chi_{D_{1}},\ldots,\chi_{D_{p}}$ are linearly dependent modulo $\mathcal{A}\mathcal{C}(G/K,\mathbb{Q})$: observe for example that $\sum_{j=1}^{p}\chi_{D_{j}}$ is the constant function $1$ almost everywhere. Thus $p-1=\dim_{\mathbb{Q}}\mathcal{S}_{2}\geq\dim_{\mathbb{Q}}\mathcal{S}_{1}=m$. Since every $D_{j}$ spans an infinite connected component of $\Gamma(G,K,S)-\mathrm{span}(F)$, one concludes that the number of ends of $\Gamma(G,K,S)$ is at least $p\geq m+1$. ∎
Corollary 7.2.
Given a compactly generated t.d.l.c. group $G$, the dimension of the $\mathbb{Q}$-vector space $\mathrm{dH}^{1}(G,\mathbb{Q}[G/K])$ is the same for all compact open subgroups $K$ of $G$.
Let $\mathcal{C}\mathcal{O}(G)$ be the poset of all compact open subgroups of $G$ ordered by reverse inclusion $\prec$. Recall that the rational discrete standard bimodule of $G$ is defined in [19, Section 4.2] as
[…]
where, for every $K\prec H$, $\eta_{K,H}\colon\mathbb{Q}[G/K]\to\mathbb{Q}[G/H]$ is the $\mathbb{Q}[G]$-module homomorphism given by $\eta_{K,H}(K)={|K:H|}^{-1}\sum_{r\in\mathcal{R}_{K,H}}rH$. Here $\mathcal{R}_{K,H}$ denotes a set of representatives of the left cosets of $H$ in $K$ that contains $1$.
Theorem 7.3.
Let $G$ be a non-compact compactly generated t.d.l.c. group. The $\mathbb{Q}$-vector space $\mathrm{dH}^{1}(G,\mathrm{\bf Bi}(G))$ has dimension $e(G)-1$ if $G$ has finitely many ends. Moreover, $G$ has infinitely many ends if and only if $\mathrm{dH}^{1}(G,\mathrm{\bf Bi}(G))$ is infinite-dimensional.
Proof.
For all compact open subgroups $H\subseteq K$ of $G$, $\eta_{K,H}\colon\mathbb{Q}[G/K]\to\mathbb{Q}[G/H]$ induces the injective map $\mathrm{dH}^{1}(\eta_{H,K})\colon\mathrm{dH}^{1}(G,\mathbb{Q}[G/K])\to\mathrm{dH}^{1}(G,\mathbb{Q}[G/H])$ (cf. [19, End of proof of Proposition 4.7]). By [15, Remark 3.4], one deduces that
[…]
where the set of all compact open subgroups $\mathcal{CO}(G)$ is ordered by reverse inclusion.
The claim then follows by Corollary 7.2 and the following two facts:
Fact 1:
[…]
Fact 2:
[…]
In order to prove Fact 2, it is sufficient to observe that, by Corollary 7.2, $\mathrm{dH}^{1}(\eta_{H,K})$ is an isomorphism for all $H$ and $K$. Finally, [45, Proposition 1.2.4(c)-(d)] yields the claim. ∎
Corollary 7.4.
Let $G$ be a compactly generated t.d.l.c. group. Then,
$e(G)=1-\dim_{\mathbb{Q}}\mathrm{dH}^{0}(G,\mathrm{\bf Bi}(G))+\dim_{\mathbb{Q}}\mathrm{dH}^{1}(G,\mathrm{\bf Bi}(G)).$
Proof.
The statement follows from the fact that $e(G)=0$ if, and only if, $G$ is compact, from Theorem 7.3 and [19, Proposition 4.3(b)]. ∎
Combining Lemma 3.2 and Corollary 7.4, one deduces what follows.
Corollary 7.5.
Let $\Delta$ be a locally finite building and let $G$ be a t.d.l.c. group acting on $\Delta$ with compact stabilisers and finitely many orbits. Then
(7.5) $e(|\Delta_{\mathrm{Dav}}|)=e(G)=1-\dim_{\mathbb{Q}}\mathrm{H}_{c}^{0}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q})+\dim_{\mathbb{Q}}\mathrm{H}_{c}^{1}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q}),$
where $e(|\Delta_{\mathrm{Dav}}|)$ is defined as in [22, Appendix G.3]. In particular, if $G$ is non-compact, then $e(G)=1+\dim_{\mathbb{Q}}\mathrm{H}_{c}^{1}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q})$.
Proof.
By Lemma 3.2 (together with the isomorphism $\mathrm{\bf Bi}(G)\to C_{c}(G,\mathbb{Q})$ in [19, Equation (4.64)]) and Corollary 7.4, one has
$e(G)=1-\dim_{\mathbb{Q}}\mathrm{H}_{c}^{0}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q})+\dim_{\mathbb{Q}}\mathrm{H}_{c}^{1}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q})$
and, if $G$ is non-compact, $e(G)=1+\dim_{\mathbb{Q}}\mathrm{H}_{c}^{1}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q})$. It suffices now to observe that $G$ acts geometrically on the proper geodesic space $|\Delta_{\mathrm{Dav}}|$ and then $e(G)=e(|\Delta_{\mathrm{Dav}}|)$ (cf. [21, Theorem 4.C.5]). ∎
Remark 7.6.
Let $G$ be a t.d.l.c. group acting on a locally finite building $\Delta$ of type $(W,S)$ with compact stabilisers and finitely many orbits. By Theorem 3.4 and [19, Proposition 3.7(a)], one has
$e(G)=0$$\Longleftrightarrow$ $G\text{ compact }\Longleftrightarrow\mathrm{cd}_{\mathbb{Q}}(G)=0$
$\Longleftrightarrow$ $\mathrm{cd}_{\mathbb{Q}}(W)=0\Longleftrightarrow W\text{ finite }\Longleftrightarrow e(W)=0.$
The following proposition provides an alternative proof of [14, Proposition 5.15], and it is based on a result of [23].
Proposition 7.7.
Let $G$ be a non-compact t.d.l.c. group acting on a locally finite building $\Delta$ of type $(W,S)$ with compact stabilisers and finitely many orbits. Then $e(W)\leq e(G)$ and
(7.6) $e(W)=1\Longleftrightarrow e(G)=1.$
Proof.
According to [22, p. 338], there is a continuous map $\rho\colon|\Delta_{\mathrm{Dav}}|\to|\Sigma(W,S)_{\mathrm{Dav}}|$ which is a retraction (i.e., it admits a right-inverse continuous map). By functoriality, $\rho$ induces the surjective linear maps
(7.7) $\rho_{k}^{*}\colon\mathrm{H}_{c}^{k}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q})\longrightarrow\mathrm{H}_{c}^{k}(|\Sigma(W,S)_{\mathrm{Dav}}|,\mathbb{Q}),\quad k\geq 0.$
By Corollary 7.5 and (7.7), one deduces that
$e(W)\leq 1+\dim_{\mathbb{Q}}\mathrm{H}_{c}^{1}(|\Sigma(W,S)_{\mathrm{{Dav}}}|,\mathbb{Q})\leq 1+\dim_{\mathbb{Q}}\mathrm{H}_{c}^{1}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q})=e(G).$
In order to prove the equivalence, one first notices that $e(G)=1+\dim_{\mathbb{Q}}\mathrm{H}_{c}^{1}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q})$ and $e(W)=1+\dim_{\mathbb{Q}}\mathrm{H}_{c}^{1}(|\Sigma(W,S)_{\mathrm{Dav}}|,\mathbb{Q})$ (cf.  Corollary 7.5 and Remark 7.6). Hence, (7.6) holds once the following is proved:
(7.8) $\mathrm{H}_{c}^{1}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q})=0\Longleftrightarrow\mathrm{H}_{c}^{1}(|\Sigma_{\mathrm{Dav}}|,\mathbb{Q})=0.$
In fact (7.8) holds for every degree $k\geq 0$. Indeed, by the main theorem of [23] one has
(7.9) $\mathrm{H}_{c}^{k}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q})\simeq\bigoplus_{T\subseteq S,\atop T\text{ spherical}}\mathrm{H}^{k}(K,K^{S-T})\otimes\widehat{A}^{T}(\Delta)$
and
(7.10) $\mathrm{H}_{c}^{k}(|\Sigma(W,S)_{\mathrm{Dav}}|,\mathbb{Q})\simeq\bigoplus_{T\subseteq S,\atop T\text{ spherical}}\mathrm{H}^{k}(K,K^{S\setminus T})\otimes\widehat{A}^{T}(\Sigma(W,S)),$
for every $k\geq 0$. The Davis chamber $K$ depends only on the Coxeter group $(W,S)$ and not on the building. By [23, Remark at p. 570 and Definition 7.4], the abelian groups $\widehat{A}^{T}(\Delta)$ and $\widehat{A}^{T}(\Sigma(W,S))$ are non-vanishing for every spherical subset $T\subseteq S$. Therefore, $\mathrm{H}_{c}^{k}(|\Delta_{\mathrm{Dav}}|,\mathbb{Q})=0$ if, and only if, $\mathrm{H}^{k}(K,K^{S-T})=0$ for every $T\subseteq S$ spherical (cf. (7.9)). By (7.10), an analogous equivalence holds for $\mathrm{H}_{c}^{k}(|\Sigma(W,S)_{\mathrm{Dav}}|,\mathbb{Q})$. ∎
Remark 7.8.
In Proposition 7.7 the equality between $e(W)$ and $e(G)$ does not hold in general. For example, consider $G=\mathrm{SL}_{2}(\mathbb{Q}_{p})$ acting on its Bruhat–Tits building $\Delta$. It is well-known that $\Delta$ is a $(p+1)$-regular tree and, by Corollary 7.5, $e(\mathrm{SL}_{2}(\mathbb{Q}_{p}))=\infty$. On the other hand, the Weyl group associated to $\Delta$ is the infinite dihedral group, which is two-ended.